### v3

# Fix-Con: Automatic Fault Localization and Repair of Deep Learning Model Conversions between Frameworks

Abstract
Converting deep learning models between frameworks is a common step to maximize model compatibility across devices and leverage optimization features that may be exclusively provided in one deep learning framework. However, this conversion process may be riddled with bugs, making the converted models either undeployable or problematic, considerably degrading their prediction correctness.
In this paper we propose an automated approach for fault localization and repair, Fix-Con, during model conversion between deep learning frameworks. Fix-Con is capable of detecting and fixing faults introduced in model input, parameters, hyperparameters, and the model graph during conversion.
Fix-Con uses a set of fault types (mined from surveying conversion issues reported in code repositories and forums) to localize potential conversion faults in the converted target model and then repair them appropriately, e.g., replacing the parameters of the target model with those from the source model. This is done iteratively for every image in the dataset, comparing output label differences between the source model and the converted target model until all differences are resolved. We evaluate the effectiveness of Fix-Con in fixing model conversion bugs of three widely used image recognition models converted across four different deep learning frameworks. Overall, Fix-Con was able to fix $462462$ out of $755755$ detected conversion faults, either completely repairing or significantly improving the performance of $1414$ out of the $1515$ erroneous conversion cases.
I Introduction
With the widespread use of deep learning (DL) in various domains, there is an inherent need for DL models to be inter-operable across DL frameworks (such as PyTorch [67], TensorFlow (TF) [64], Keras [54]) to maximize re-usability of models across frameworks. Conversion of DL models between frameworks is facilitated by a plethora of automated conversion tools such as tf2onnx [44], onnx2keras [45], onnx2torch [46], MMdnn [61], among others. However, the conversion process may be riddled with bugs [63, 62, 43], making the converted models undeployable, perform poorly in terms of output label correctness when changing from one framework to another, run slowly, or face challenges in robust deployment [66, 52, 53].
Louloudakis et al. [63] provided empirical evidence of model conversion errors and resulting output label errors in four DL frameworks. In this paper, we target the conversion errors they observed for fault localization and repair. They considered three image recognition models pre-trained on ImageNet [55] that were treated as Source models. They then converted each Source model to use a different DL framework, referred to as Target. Figure 1 from [63] shows the proportion of output label dissimilarities between (Source, Target) pairs across all images in the test dataset. The conversion process failed to execute in $1111$ out of the $3636$ conversions, seen as empty gray cells in Figure 1. For the remaining conversions, they found varying levels of image label discrepancies between Target and Source models from 0 to 100%, with $1515$ cases having non-zero percentage of output label dissimilarity. This study underscores the error-prone nature of the conversion process, highlighting the necessity for a technique to localize and repair faults introduced by DL framework converters.
Figure 1: Pairwise comparison of output labels between Source and converted Target models (from [63]).
To address this problem, in this paper we propose Fix-Con, an automated approach for fault localization and repair of erroneous model conversions between DL frameworks. We focus on the error cases in Figure 1 with non-zero output label differences. The gray boxes in Figure 1, where the conversion tools crashed, have been explored by existing tools like MMdnn [61] and complements the capability we provide in diagnosing and repairing conversions with output label differences.
Fix-Con is capable of detecting and fixing faults introduced in model layer weights, biases, hyperparameters, and the model computation graph during conversion. It compares the Source model (in one DL framework) against the converted Target model (in a different framework) using an input dataset. For images with label discrepancies between Source and Target, Fix-Con assesses the difference between the models with respect to their parameters. Then it uses a set of strategies to mitigate conversion errors such as the replacement of Target model parameters with those from Source.
Our approach for Fix-Con focuses on the Target model rather than the converter tool itself for a variety of reasons. First, errors observed in conversions may be related to problematic configurations (e.g., input preprocessing) that are not related to the conversion tool. Second, the source code of the conversion tool may be unavailable or proprietary, and therefore not accessible for direct repair. Third, the conversion process might involve manual intervention or conversion (e.g., [25, 6]), which in turn can introduce errors. Fourth, our methodology is not tied to any specific conversion tool/version, and can be applied to any model conversion between source and target settings, including intermediate representations like ONNX [47]. It is worth noting that conversion tools are constantly evolving, thus having a localize/repair approach independent of the tool source code accessibility and evolution would prove helpful and more dependable for AI model users. As an aside, the fault localization and repair techniques we propose can also be adopted by conversion tool developers to address faults associated with the tools themselves.
We evaluate the effectiveness of Fix-Con in fixing the model conversion errors reported by Louloudakis et al [63]. Fix-Con was able to effectively localize and repair $1414$ of the $1515$ erroneous model conversions that had non-zero output label dissimilarities.
In summary, we make the following contributions:
  1. 1.
    A novel method for localizing faults in DL model conversion between a range of DL frameworks.
  2. 2.
    An automated fault repair framework, Fix-Con, that can fix converted Target models to match output labels with that of the Source model.
  3. 3.
    A comprehensive empirical evaluation of the effectiveness of Fix-Con in fault localization and repair.
II Background
We provide a brief background on deep learning frameworks, conversions between them, the intermediate ONNX format and the compiler framework Apache TVM that we use for implementation and analysis with Fix-Con.
II-A Deep Learning Frameworks
Deep Learning Frameworks provide utilities to machine learning engineers and developers in order to build, train, deploy and optimize Deep Neural Networks (DNNs). We focus on four actively maintained and widely-used DL frameworks in our experiments: Keras [54], PyTorch [67], TensorFlow (TF) [64], and TensorFlow Lite (TFLite - a framework focused on edge devices) [64].
II-B Framework Conversions
DNN models are converted from one DL framework to another, for two primary reasons: (1) portability - so that a model can be deployed on devices of different computational capabilities; and (2) supported features and/or operations, such as extended optimization support or better profiling options. The conversion process is conducted by providing a pre-trained model built on one DL framework (e.g., PyTorch). We call this model Source. Then, the conversion tool will use Source in order to generate the same model, but built in a new DL framework (e.g., Keras). We call this model Target. There is a wide variety of open-source tools for this purpose [44, 45, 46, 61], and the conversion procedure is performed automatically. However, the conversion process might involve multiple converters and might also require the model to be built using an intermediate representation (e.g., ONNX [47]). For instance, the conversion of a model from PyTorch to Keras might require a two-step conversion: (1) PyTorch to ONNX, using the native converter of PyTorch to generate ONNX; and (2), the onnx2keras [45] converter, in order to convert the ONNX intermediate representation to Keras.
II-C Open Neural Network Exchange (ONNX)
ONNX [47] is an open standard for machine learning interoperability that allows the representation of DNNs using a common set of operators and file formats to enable interoperability with a variety of tools, and systems. ONNX provides a powerful API in order to analyze model graph, hyperparameters, and parameters. The model is represented by a graph of operations (or nodes), each containing a set of properties. These operations are executed sequentially, propagating data from one layer to its subsequent ones. Throughout the years, ONNX evolves producing newer versions of the standard, modified from the past. Each standard version is attributed an opset number. Due to the popularity of ONNX and the versatility of its API, we perform our automatic fault localization and repair in this representation level - retrieving ONNX versions of the Source and Target model, analyzing them and repairing them. For that purpose, when we refer to Source and Target  we implicitly refer to the ONNX variants of the models under test.
II-D Apache TVM
Apache TVM [51] is an end-to-end machine learning compiler framework for CPUs, GPUs, and specialized accelerators. It is actively used and supported by a wide community of developers and researchers worldwide. TVM allows compilation, optimization and deployment of DNNs in a variety of hardware acceleration devices. It also features a very powerful debugger, which allows to export metadata regarding the model inference, such as layer activations, execution times per-model layer, as well as metadata about model structure and parameters, each exported in separate metadata files, while also providing a high-level API for their analysis. We utilize the TVM Debugger in order to perform Layer Activation Analysis, as described in Section IV-A4.
III Related Work
In terms of model conversions, limited research is conducted in the literature. Openja et al. [65] published a study that emphasizes the challenges of the conversion process. Louloudakis et al. [63, 62] explored the effects of the computational environment parameters on image recognition models, including deep learning (DL) framework conversions. Liu et al. [61] propose MMdnn, a framework to automate the framework conversion process and mitigate a number of conversion challenges that result in conversion failures, such as unavailable operators, inconsistent tensor layouts, unsupported padding and incompatible argument types. MMdnn, however, does not consider faults related to configuration, such as input preprocessing, or faults related to additional important aspects such as layer parameters and computation graph that can also result in output incompatibilities during conversion. This is addressed by Fix-Con. It is also worth noting that MMdnn is currently in a deprecated state, with its last update on August 2020. In addition, our efforts with testing MMdnn was unsuccessful as it was built with older, deprecated versions of DL frameworks (e.g., PyTorch v0.4.0, while the current version is v2.2.0).
In an attempt to compare MMdnn with our work, we identified a potential overlap of only $1515$ potential faults out of the $462462$ we repaired with Fix-Con, based on the results reported in [61]. Finally, MMdnn supports conversions over a restricted set of DL frameworks (primarily CNTK, CoreML, Keras, MXNet, PyTorch, TensorFlow, Caffe, and DarkNet) that is not currently maintained or updated while our approach is generalizable to model conversions between any pair of DL frameworks.
Automatic fault localization of DNNs is also explored in the literature DeepFault [56] focuses on a suspiciousness-oriented spectrum analysis algorithm to identify problematic parts of the model that can be responsible for faults, and also proposes a method for adversarial input generation. DeepLocalize [74] converts the model into an imperative representation and then performs dynamic analysis on its execution traces. In the context of fault localization of DL frameworks, CRADLE [68] performs an execution analysis on the model graph in order to detect faults, while LEMON [73] utilizes the metrics used by CRADLE and applies mutation testing in order to identify issues in the model. However, none of the aforementioned approaches focus on faults encountered during model conversions.
To achieve model conversions between DL frameworks, a variety of popular tools exist, such as tf2onnx [44], onnx2keras [45], onnx2torch [46], tflite2onnx [42], and native DL framework converters such as TFLiteConverter from TFLite [64], and ONNX converter from PyTorch [67]. However, these tools come with no guarantees, and can potentially introduce faults in the converted models.
IV Methodology
Figure 2: Fault localization and repair pipeline of Fix-Con handling 6 fault types, three within Input-based category and three within Layer-based category.
Fix-Con first takes as input the dataset, Source and Target models. It then localizes and repairs conversion errors that resulted in output label discrepancies between Source and Target. To compare the Source and Target model architectures, configuration and parameters we use their corresponding representation in the ONNX format which is a popular intermediate representation. We then import the Source-ONNX and Target-ONNX models to the TVM compiler framework to facilitate comparison. Fix-Con comprises two main components: 1) A Fault Localization component that identifies locations, parameters and properties in the Target model that may be potentially faulty; and 2) a Fault Repair component with a defined set of strategies to fix the faulty locations and aspects identified through fault localization. The full fault localization and repair pipeline of Fix-Con is presented in Figure 2.
IV-A Fault Localization
We surveyed $4040$ posts and issues from StackOverflow, GitHub, and forums such as TensorFlow and NVidia discussions, where model conversion errors were raised, to understand the possible locations of errors. Table I shows six different conversion fault types identified in these issues that were related to 1) Preprocessing (PP), (2) Input Dimensions (ID), (3) Tensor Shape and Structure (TSS), (4) Weights & Biases (WB), (5) Layer Hyperparameters (LH), and (6) Computation Graph (CG). We also discovered many of these fault types examining the conversion errors reported in [63].
TABLE I: Model conversion faults from StackOverflow, GitHub issues and other forums. Faults are classified as input-based and layer-based.
# Class Fault Sources
1 Input $P$ [3, 19, 10, 39, 7, 30]
2 Input ID [5, 36, 37, 32, 13, 20, 23, 26, 28]
3 Input TSS [11, 14, 21, 27, 31, 41]
4 Layer WB [8, 9, 12, 16, 2, 29, 34, 1, 15, 35, 33, 40]
5 Layer LH [17, 18, 50]
6 Layer CG [4, 38, 22, 24]
As seen in Figure 2, our fault localization approach starts by localizing the first three fault types: PP, ID, and TSS. These three fault types pertain to the input, stemming from differences between Source and Target with respect to the preprocessing used, or changes in input dimensions or operations on the input. We refer to these fault types as Input-based. After attempting to localize input-based fault types, our approach proceeds with the remaining three fault types in Table I that belong to the Layer-based category, as it stems from differences in layer hyperparameters, weights and biases, and differences in computation graph.
Comparisons for Input-Based Fault Types
We examine and contrast input-related settings between the Source and Target models in the fault types below.
IV-A1 Preprocessing (PP)
For some instances, differences in preprocessing settings between Source and Target can result in output label differences. We explore the effect of different preprocessing configurations for Targetusing the default preprocessing setting from the official repositories of the DL frameworks both Source and Target. For example, in the case of converting MobileNetV2 from PyTorch to Keras, Fix-Con examines the performance of the converted Target, using both official PyTorch and Keras preprocessing settings for MobileNetV2.
IV-A2 Input Dimensions (ID)
Islam et al. [60] identified that changes in layer input dimensions can affect the performance of a DNN model. We also empirically identified that model conversion tools accept misconfigurations of Target model input dimensions without notifying the user of the mismatch between Source and Target input dimensions that may potentially result in output label discrepancies. We explore problems arising from such misconfigurations in this fault type. For example, we inspected the conversion of ResNet101 from PyTorch to TFLite, modifying the input dimensions from (1, 3, 224, 224) used in Source to (1, 3, 299, 299) used in Target. The conversion process completed without errors and warnings. However, the model presented a significant output label discrepancy between Source and Target (37373737%). In our fault localization approach, we check if input dimensions between Source and Target match, and report any mismatches to the fault repair component.
IV-A3 Tensor Shape and Structure (TSS)
Another potential error identified by Islam et al. [60] relates to tensor shapes and changes in model structure. We consider cases where an erroneous tensor shape can affect the correctness of the model conversion where Source and Target models have differences in output labels. Such a scenario may be observed when a transpose operation is introduced by a converter tool to modify the model input. This is typically manifested as a Transpose node right after the input node in the computation graph. In an erroneous conversion scenario, the Transpose node might generate a tensor with correct shape, but incorrect structure. To illustrate this, consider the example of a Source input with the shape (1, 224, 224, 3). The Target model on the other hand has a shape of (1, 3, 224, 224) after applying a Transpose node that modifies the input of Target, by either applying Transpose(tensor, [0, 2, 3, 1]), or Transpose(tensor, [0, 3, 2, 1]), depending on the implementation. The modified shape is used in subsequent nodes in the computation graph. Owing to the ambiguity in dimensions where both transpose operations result in a tensor with correct shape expected by the subsequent layers in the Target model, it can result in model conversion errors as one of them has an incorrect structure. To localize faults like this, Fix-Con checks if there is a difference in model input shape between Source and Target. For cases presenting differences, Fix-Con further examines if any of the input-related subsequent layers contains a shape transformation operation (like Transpose) and considers such scenarios as potential sources of faults, conservatively. Potential faults related to tensor structure may also happen when a layer tensor in the Target model is incorrectly transformed by the converter and is followed by a flattening or a reshaping operation that loses the original structure information, making the difference with Source model non-obvious. In particular, if a tensor has a different shape between Source and Target (i.e., due to erroneous conversion) but the total number of elements is the same across tensors, a Flatten or a Reshape operation on such a tensor will make it impossible to detect the mismatch between Source and Target, as the structure information is lost. However, the order of elements between the corresponding Source and Target tensors will be different after the Flatten or a Reshape operation, which may result in output label discrepancies. For instance, if a layer has shape (1, 3, 32, 32) on Source  but was converted to (1, 32, 32, 3) on Target, and is followed by a Flatten node in both models, the result will be tensors with (3072) elements in both models. We localize this fault by examining the tensor shape of the dominator nodes preceding any Flatten or Reshape nodes, and check if it matches with the corresponding nodes in the Source model.
Comparisons for Layer-Based Fault Types
For this category of faults, we examine and compare layer activations between Source and Target models. For layers where differences in activations are outside of an expected range of differences (as described in Section IV-A4), we mark those layers as suspicious and rank them so that we examine the most suspicious layer first for fault types CG, LH, and WB. We then iteratively proceed down this ranked list of suspicious layers for inspecting the three layer-based fault types, CG, LH and WB, discussed below. We start by describing our approach for producing a ranked list of suspicious layers in layer activation analysis followed by our fault localization approach for CG, LH, and WB with the ranked suspicious layers. An example instance of the differences between Source and Target in weights, biases, and hyperparameters that Fix-Con attempts to localize as potential faults is shown in Figure 3. We find, for a core computational layer under inspection, that the weights present small differences between Source and Target. Additionally, the pads hyperparameter that is present in Source, is missing on Target in Figure 3.
IV-A4 Layer Activation Analysis
We compare the activations for each convolutional layer between Source and Target models, as a four-step process. (1) We extract two equally sized sets of images from the test dataset: one for similar images (images that produce the same top label prediction across Source and Target), and one for dissimilar images (images with different labels between Source and Target). We then extract the layer activation tensors of both sets by loading the Source and Target ONNX models to TVM and performing inference using the TVM Debugger [49]. The TVM debugger generates the complete model graph and parameters at compile time, while also gathering layer activations for each input. Using these activations and the Source ONNX model, we then utilize the LAYER_ANALYSIS function, shown in Algorithm 1. This function takes as input the original source ONNX model, as well as the activations for the sets of similar and dissimilar images (sim_acts, dis_acts). It also utilizes the List module, responsible for list length check and sorting operations, and the Kruskal-Wallis parametric test (KWTest). (2) To identify when activation mismatches between Source and Target indicate a potential problem, we first extract the differences in activations between Source and Target, per element within each layer, for the set of similar images and store these differences in a set, sim_diff in line $88$ of Algorithm 1. We treat differences in activations between Source and Target for these similar images as acceptable differences and we use it to estimate the expected difference in distribution of activations for correct conversions. (3) Next, for the set of dissimilar images, we compute the set of dissimilar differences, diss_diff in line $1010$, per tensor element per layer. We then check if the difference in activations per tensor element per layer belongs to the corresponding expected distribution of activation differences (from the previous step). To do this, we use the Kruskal-Wallis non-parametric statistical test (line $1111$) [57] [^1] The distribution of activation differences for correct conversion instances did not follow a normal distribution, so we chose the non-parametric Kruskal-Wallis test. to determine if sim_diff and diss_diff belong to the same distribution (at 5% significance level) (line $1515$). (4) From step 3, we obtain the tensor elements whose difference in activations for sim_diff and diss_diff did not belong to the same distribution, for each layer between Source and Target. These are considered as potentially problematic elements that may result in model conversion errors. We count the number of problematic tensor elements for each model layer (line $1616$) and rank the layers in descending order of number of problematic elements (line $1818$). This ranking of suspicious layers is the basis for fault localization and repair of fault types CG, LH, and WB. We only consider convolutional layers as these are the core computational layers for a Convolutional Neural Network (CNN). However, this can be modified to consider other layer types as well.
Algorithm 1 Layer Analysis in Fix-Con
1:List, KWTest, model, sim_acts, diss_acts
2:procedure layer_analysis(model, sim_acts, diss_acts)
3:    for layer in model.layers do
4:         layer.prob_elems $\leftarrow$ 0
5:         for element in layer.elements do
6:             sim_diff, diss_diff $\leftarrow$ [], []
7:             // Gather element-wise activations.
8:             for sim in sim_acts do
9:                 sim_diff $\leftarrow$ sim_diff $\cup$ sim[element]
10:             for diss in diss_acts do
11:                 diss_diff $\leftarrow$ diss_diff $\cup$ diss[element]
12:             kw_out $\leftarrow$ KWTest(sim_diff, diss_diff)
13:             // If the null hypothesis of
14:             // same distribution is rejected,
15:             // then the element counts as problematic.
16:             if kw_out.p_value <<< 0.05 then
17:                 layer.prob_elems+++++ +
18:    // Sort layers in descending order of problematic elements.
19:    return sort(model.layers,
              (l1, l2) $\Rightarrow$ l1.prob_elems -- l2.prob_elems)
IV-A5 Weights and Biases (WB)
We extract the weight and bias tensors for each of the convolutional layers. Starting from the most suspicious layer in the ranked list, we compare the corresponding Source and Target tensors element-wise for each of the layers. We flatten the tensors to facilitate comparison. For correct model conversions, we expect the weights and biases for the Source and Target model layers to match exactly. Values for model weights and biases are only affected during model training or optimization, neither of which is part of model conversion. In addition to convolutions, we also perform comparison of weights and biases in neighboring layers, e.g., batch normalization nodes preceding and/or succeeding the convolutional nodes). Consequently, differences identified during the comparison of weights and bias tensors for layers across Source and Target point to potential problems in those layers. We output all the layers that have differences identified in weights and bias tensors between Source and Target  and pass it to the repair component.
IV-A6 Layer Hyperparameters (LH)
Incorrectly converted hyperparameters are another potential source of error. For example, with a convolutional layer, we expect that the padding, strides, dilation, and other configurations to remain unchanged during model conversion. Starting from the most suspicious layer in the ranked list, we compare layer hyperparameters and attributes – (pads, strides, kernel_shape, dilations, epsilon, min, max, axis) between Source and Target. If we find any differences, we mark them as potential faults. In particular, we observe $33$ types of differences: (1) a hyperparamater is present in Source but not in Target; (2) a hyperparameter is present in Target but not in Source (e.g., added by the converter tool); and (3) a hyperparameter is present both in Source and Targetbut their values are different.
IV-A7 Computation Graph (CG)
In our experiments, we found that output label discrepancies can sometimes arise from differences in model graph between Source and Target models. We examine differences in graph nodes. To make the analysis tractable, we start from the most suspicious layer (based on the ranking from layer activation analysis) in Target and extract the subgraph between this layer and its dominator in the Target model graph. We do the same with the Source graph starting from the corresponding Source layer node and its corresponding dominator. We perform a depth first traversal of the two subgraphs and compare their nodes. Any differences are marked as a potential fault. We repeat this for any other suspicious layers in the ranked list.
Figure 3: Indicative example of differences in layer weights and hyperparameters, introduced in the model conversion process.
IV-B Fault Repair
We implemented a number of strategies to repair the potential faults identified by the fault localization component. We discuss the strategies to fix each of the issues identified in fault localization below. As seen in Figure 2, for layer-based repairs to CG, LH, and WB, the fault repair component will start from the most suspicious layer in the layer ranking provided by the layer activation analysis (within fault localization) and apply fixes sequentially and iteratively to layers in that ranking. Each of the fixes to Target produces an updated version. The updated Target will only be accepted if the output label discrepancy with respect to Source reduces after the repair (compared to previous target version), and this is then used as the current Target for the next repair strategy or suspicious layer in the list. On the other hand, for PP, ID, TSS, the repair actions are layer-independent, and do not need to consider the layer ranking. The repair strategies are applied sequentially with PP applied first, followed by ID and then TSS. As with layer-based strategy, the fix is only accepted if the output label discrepancy is reduced. Finally, both input-based and layer-based strategies are applied to every image in the dataset that had different labels between Source and Target until they are all fixed or the algorithm times out.
Input-based Repair Strategies
IV-B1 Preprocessing Differences (PP) Repair
We experiment running the Target model with both the Target and Source preprocessing settings to check if the label discrepancies are resolved.
IV-B2 Input Dimensions (ID) Repair
Following fault localization, if a difference is detected between Source and Target input dimensions, Fix-Con automatically fixes Target to match with the input dimensions from Source. We observed four erroneous model conversions in our experiment with incorrect input dimensions, with two of them requiring further actions in model nodes for tensor shape and structure (discussed below).
IV-B3 Tensor Shape and Structure (TSS) Repair
Following fault localization, as described in Section IV-A3, if the input is different between Target and Source in a problematic conversion, and input is succeeded by a Transpose node, then Fix-Con attempts to apply a fix by removing that node, in addition to adjusting the Target input to match Source. Furthermore, if Fix-Con detects any nodes that are related to input processing and normalization (e.g., InceptionV3 architecture contains Gather, Unsqueeze, Add, and Mul nodes before the added Transpose), then Fix-Con adjusts the attributes of these nodes in order to be compatible to the Source input dimension. In cases where Fix-Con detects the presence of Flatten or Reshape nodes that alter the shape of tensors, then it attempts to repair this by inserting a Transpose node right before Flatten or Reshape to match tensor shape with Source.
Layer-based Repair Strategies
IV-B4 Weights and Biases (WB) Repair
If a mismatch is found between any weight or bias tensor across model layers by the fault localization component, then Fix-Con replaces the values from the corresponding tensor in Source to Target. As mentioned earlier, Fix-Con performs the repair starting from the most suspicious layer in the ranked list.
IV-B5 Layer Hyperparameters (LH) Repair
Upon marking a layer hyperparamter case as potentially problematic, as described in Section IV-A6, Fix-Con attempts to fix Target by applying the corresponding setting of Source. In particular, if (1) a node is present in Source but is missing from the computation graph of Target, Fix-Con will update the corresponding node including the Source hyperparameter to Target. If (2) a hyperparameter is added in Target but is missing from Source, Fix-Con will remove it from the target node. And finally, if (3) there is a value mismatch between Source and Target for the hyperparameter, Fix-Con will update Target in accordance to the source value.
IV-B6 Computation Graph (CG) Repair
In accordance to Section IV-A7, if a difference in subgraphs is detected between Source and Target, Fix-Con attempts to repair it by replacing the Target subgraph nodes with Source subgraph nodes, preserving all the node properties and ensuring all weights, biases and hyperparameters are correctly updated in the Target graph.
Regarding the order in which our repair strategies are applied, we first applied the ones that were input-based, namely PP, ID, and TSS. We do PP first as changes to preprocessing may fix errors observed downstream. We then examine ID, with TSS being an additional restriction to ID mismatches. We then apply repairs that are “layer-based” – WB, LH, and CG. Among the ones with layer-based effects, we first perform repairs related to model weights, biases and hyperparameters, and then those related to computation graph modifications. Note that some graph transformations may also be related to hyperparameter changes, e.g., removing a Target Pad node to adapt to Source structure requires the addition of a padding hyperparameter to the next convolutional layer.
We implemented Fix-Con in Python. We used Apache TVM Debugger for layer activations analysis. The debugger allows the extraction of model graph structure and parameters in metadata files, and a log of the layer activations throughout the model structure. In order to implement the repair strategies, we used the ONNX API to modify Target.
V Experiments
We consider three widely used image recognition models of various sizes: MobileNetV2 [71], ResNet101V2 [58, 59], and InceptionV3 [72]. We use models pre-trained on ImageNet [55] using native model definitions and pre-trained parameters/weights sourced from four different DL frameworks’ repositories: Keras [54], PyTorch [67], TensorFlow(TF) [64], and TFLite [64]. For our experiments, we used two datasets: the ILSVRC2017 [70] object detection test dataset consisting of $55005500$ images, and the ImageNetV2 dataset [69] consisting of 10,0001000010,00010 , 000 images. We used the ILSVRC2017 dataset for detecting and repairing errors in model conversion. We then assessed the correctness of the repaired models on ImageNetV2, utilizing it as an independent dataset. We used the problematic conversion cases identified in [63] as our candidates for fault localization and repair. We focused on cases where the conversion process introduced label discrepancies between Source and Target, and there were $1515$ such erroneous conversions, as seen in Figure 1.
We evaluate the following research questions in our experiments:
RQ1:
How effective is Fix-Con at localizing faults in the $1515$ erroneous model conversions?
RQ2:
How effective is Fix-Con at repairing the faults in the $1515$ erroneous model conversions? We evaluate which of the repair strategies are used most frequently in model fixes.
To check the robustness of Fix-Con with respect to devices, we ran our experiments on two devices of varying capabilities, an Intel-based server featuring a high-end Nvidia Tesla K40c (GK11BGL) GPU, and a Laptop featuring a low-end Intel(R) GEN9 HD Graphics NEO. Our experiment results remained the same on both devices.
V-A Fault Analysis Setup
Within layer activation analysis, discussed in fault localization for layer-based types, we determine the expected distribution of differences in activations per layer between Source and Target for images where the output label is the same using a random sample of up to $100100$ images that produced similar results. We then compare the activations for dissimilar images against this expected distribution to come up with a ranking of suspicious layers. We repeat this process $33$ times with three different random samples of similar images.
V-B Fault Repair Setup
Once we extracted the order of layers from the fault analysis, we utilized the fault repair component of Fix-Con, applying the strategies in a greedy approach, as described in Section IV-B. We set a time limit of 2 hours for each conversion case. We also set a limit of three unsuccessful repair iterations, when repairs did not reduce output label discrepancy, as a stopping condition. This is to help make the experiments tractable. These settings can be customized as needed in our tool.
VI Results
Overall, Fix-Con was able to localize $755755$ potential faults in the structure and properties between Source and Target variants across all $1515$ erroneous conversion cases. An overview of the potential faults distribution can be found in Table II. From these $755755$ cases Fix-Con identified as bugs, it repaired $462462$ of them — completely repairing $1212$ out of $1515$ problematic model conversions, while significantly improving the performance of $22$ more, with the most distinctive case being improved from 48.948.948.948.9% output label discrepancy to 0.330.330.330.33%. Our system was unable to reduce the percentage further, as none of our strategies affected the model beyond this point.
VI-A Fault Localization
The faults localized were distributed across all six types within input-based and layer-based fault categories. Of the $755755$ faults localized, $462462$ of them (61.161.161.161.1% of total cases) resulted in reducing output label discrepancy between Source and Targetwhen repaired. The remaining localized faults did not result in any improvement after repair and are treated as false positives detected by the fault localization component. We demonstrate the distribution of faults detected and fixed across the different types in Table II and discuss each type in the context of fault localization next.
TABLE II: Number of faults localized versus those repaired (#Localisations/#Repairs) per model conversion case.
#Localisations/#Repairs
Input-Based Layer-Based
# MobileNetV2 $P$ ID TSS WB LH CG
1 Keras-to-Torch 1/1 0/0 0/0 0/0 6/0 0/0
2 Keras-to-TFLite 1/1 0/0 0/0 96/96 0/0 0/0
3 Torch-to-TF 1/1 0/0 0/0 0/0 6/0 0/0
4 Torch-to-Keras 1/1 1/1 1/1 0/0 39/0 0/0
5 TF-to-TFLite 0/0 0/0 0/0 68/66 0/0 0/0
6 TF-to-Keras 1/1 0/0 0/0 0/0 40/0 4/0
ResNet101
7 Keras-to-Torch 1/1 0/0 0/0 0/0 72/0 2/0
8 Keras-to-TFLite 0/0 0/0 0/0 180/174 33/0 37/0
9 Torch-to-TFLite 1/1 1/1 0/0 0/0 0/0 0/0
10 Torch-to-TF 1/1 0/0 0/0 0/0 0/0 0/0
11 TFLite-to-Keras 0/0 0/0 0/0 0/0 4/4 3/3
12 TF-to-TFLite 0/0 0/0 0/0 0/0 0/0 0/0
13 TF-to-Keras 0/0 1/1 1/1 0/0 1/1 1/1
InceptionV3
14 Torch-to-Keras 1/1 1/1 1/1 0/0 54/7 0/0
15 TF-to-TFLite 0/0 0/0 0/0 94/94 0/0 0/0
Total 9/9 4/4 3/3 438/430 254/12 47/4
Preprocessing (PP): We detected $99$ cases in which the choice of preprocessing between Source and Target configurations was a decisive factor for the model performance, while $55$ cases presented the same results across Source and Target configurations. All models presented severe performance degradation in output correctness if no preprocessing option was selected.
Input Dimensions (ID): Fix-Con identified four model cases where the input dimensions were different between Source and Target. We identified that input dimension errors had two primary causes: (1) erroneous configuration from the user, which neither the model converters nor the model itself raised as warnings or errors; (2) implicit changes in model input due to the converter implementation.
We observed two cases, $99$ and $1313$ in Table II, where ID fault is a result of the first cause. Regarding the second cause, we detected two cases where the converter changed the model input. In addition to ID faults, we also detected differences in tensor shape and structure that we discuss as part of the next category.
Tensor Shape and Structure (TSS): Fix-Con detected three cases with potential faults in the TSS category. The model converter introduced a Transpose node in these two cases to address the difference in ID so that subsequent nodes use inputs with the correct dimension. However, in these two cases getting to the correct input dimension with the Transpose node was not adequate. This occurred due to similar sized dimensions within tensors, like the third and fourth dimension within an example tensor with shape (1, 3, 224, 224), where it is possible to incorrectly transpose dimensions of the same size that result in a different structure from what is intended.
In addition, in some model architectures (e.g., InceptionV3) the Transpose node is placed after a series of nodes that normalize data based on a specific input dimension. For instance, in InceptionV3, the Transpose node is placed after a series of Gather, Unsqueeze, Mul, and Add nodes. Fix-Con detects these differences across model graph and reports them during fault localization.
Weights and Biases (WB): We observed $438438$ cases of differences in tensor values between weights and biases across Source and Target, with ResNet101 converted from Keras to TFLite containing $177177$ of them.
Layer Hyperparameters (LH): Fix-Con detected $254254$ potential faults across $1212$ model conversion cases, with most of them concerning padding hyperparameters. Fix-Con also identified some cases where a potential hyperparameter fault was associated with a potential computation graph (CG) fault - for instance, specific convolutional layers presented potential faults related with padding hyperparameters and Pad nodes that directly preceded these convolutional layers.
Computation Graph (CG): Our system identified $4747$ cases across $44$ conversion cases in which parts of the model computation graph presented discrepancies across Source and Target variants. All the cases observed were associated primarily with three operations: (1) setting a padding around a tensor, by utilizing Pad nodes; (2) performing batch normalization to a tensor by applying an element-by-element constant addition, and scaling using a BatchNorm node; and (3) redundant output nodes erroneously placed by the converter. From those $4747$ cases, Fix-Con successfully identified $44$ of those cases as actual bugs and repaired them, playing a crucial role for the repair of cases #$1111$ and $1313$.
VI-B Fault Repair
[…]
PP: As shown in Table III, Fix-Con was able to repair cases #$11$, $33$, $66$, $77$, $99$ and $1010$, by determining and setting the correct preprocessing configuration. Given that the preprocessing is not a standard procedure across models, we identified that for these cases using the Source preprocessing instead of the one officially provided for the Target library fixed the result completely. Cases #$22$, $44$, $99$, and $1414$ were also affected partially by preprocessing selection, however additional steps were needed in order to repair them. The additional steps are explained in more detail in the categories below.
ID: Among the $44$ faults localized as ID, Fix-Con repaired one case completely by solely adjusting the input dimensions of Target to match Source. For the remaining three cases, additional actions had to be performed, involving repair tensor shape and structure, as well as computation graph repair, as described in the next categories.
TSS: Fix-Con was able to repair three cases related to problematic tensor shape and structure. In particular, cases $44$, $1313$, and $1414$ were related to problematic layers right after input. All three cases had different input dimensions between Source and Target, and the conversion tool had inserted additional Transpose nodes in order to adjust the input tensor shape so that it would be aligned with subsequent nodes. However, the transformation was done improperly, leading to severe performance degradation of up to 68.2168.2168.2168.21%. We identified all cases to be related to the onnxmltools [48] converter (part of the native ONNX codebase), when converting Keras to ONNX. Fix-Con automatically identified the presence of a redundant Transpose node following the input node, and repaired it by (1) adjusting the Target model input dimensions so that they correspond to those of Source, (2) altering its tensor transposition order (permutation) property to align with the new input dimensions, and (3) updating the properties of nodes connected to Transpose to ensure proper tensor handling, reflecting the change in Transpose. This modification in particular was vital to case #$1414$ where Fix-Con detected additional nodes (Gather, Unsqueeze, Add and Mul) preceding Transpose that apply input normalization. As part of the repair process, Fix-Con updated the axes properties of Gather and Unsqueeze nodes to align with the updated input dimension and the Transpose permutation in Target.
WB: Cases #$22$, $55$, $88$, and $1515$ were related to parameter precision discrepancies between Source and Target. Although the discrepancies were negligible, the model behavior deviated across the cases from 3.123.123.123.12-9.79.79.79.7%. Fix-Con performed multiple iterations and managed to completely fix cases #$55$ and $1515$ while also significantly improving cases $22$ and $88$ by replacing the weights of Target with those of Sourcewhere differences were detected. We identified all cases to be related to the behavior and the setup of TFLiteConverter. The discrepancy after each repair step is depicted in Figure 4. At each repair step, Fix-Con adjusts the weights & biases of a suspicious layer in Target with those from Source, in the order determined by the Layer Activation Analysis (Section IV-A4), and evaluates the effects of the change by comparing the predictions of Source and Target.
LH: From the $254254$ potential faults Fix-Con localized, only $1212$ of them were actively affecting the model. These hyperparameters were associated with the pads hyperparameter of Conv and MaxPool nodes (cases #$1111$ & $1313$), as well as with the axes of Gather and Unsqueeze nodes responsible for input processing (case #$1414$). Repairing Target entailed adjusting their values with those from Source.
CG: Fix-Con identified $4747$ cases of graph-related potential faults. Of these, $44$ cases resulted in fault repairs that improved model performance (reduced output label discrepancy). The repairs played a crucial role in conversion cases #$1111$ and #$1313$ that initially presented discrepancies of 100100100100% and 35353535%, respectively. Our tool identified that, for case #$1111$, the converter had erroneously introduced multiple output nodes. In addition, three problematic padding nodes were found to affect the behavior of the models. Fix-Con identified such faults and completely repaired the models by (1) removing any additional output nodes added to the model by the converter, and (2) repairing problematic padding nodes.
For all the false positive cases observed in LH and CG, we performed manual inspection and observed that the converters performed semantically equivalent computation graph conversions. For instance, in the erroneous conversion presenting the largest number of CG fault localizations – ResNet101, Keras to TFLite (case #$88$) – Source contained Pad nodes which were absent in Target. However, Target contained a pads hyperparameter in the next convolutional layer that had the same effect. Also, BatchNorm nodes of Source were replaced by the converter with Mul and Add nodes in Target  effectively applying the normalization in a two-step semantically equivalent manner. Fix-Con attempted to repair these scenarios by replacing them with the corresponding Source subgraph but these fixes had no effect on the number of output label discrepancies encountered.
In terms of the number of iterations needed during the repair process, input-based strategies are only applied once. Layer-based strategies typically require multiple iterations. Four erroneous model conversions that required multiple repair iterations related to WB strategy repair are presented in Figure 4, ranging from $22$ to $77$ repair cycles, using the ILSVRC2017 dataset.
Figure 4: Model conversion cases that required an iterative weights and biases repair strategy with the percentage output label dissimilarity shown after each repair cycle (following preprocessing repair).
VI-C Fault Localization/Repair Effectiveness
We consider effectiveness as the number of localized faults that could be repaired to improve Target prediction accuracy towards matching this of Source. We go through each fault category to determine the effectiveness of the fault localization component, as seen in the Total row in Table II. We find that Fix-Con had an 100100100100% efficacy rate (all potential fault occurrences detected could be repaired) for preprocessing (PP: 9/9999/99 / 9), input dimensions (ID: 4/4444/44 / 4) and tensor structure (TSS: 2/2222/22 / 2), as well as 98.198.198.198.1% efficacy for weights and biases (WB: 430/438430438430/438430 / 438). The weights and biases category had the most faults detected and repaired. Fix-Con had much less success on cases related to the computation graph (CG), with $44$ of the $4747$ potential faults detected being repaired (8.58.58.58.5%) while other cases when repaired did not result in reduction of output label discrepancy. The CG fault repairs, although few in number, were instrumental in completely fixing conversion cases #$1111$ and $1313$. The effectiveness for LH category was low (4444%), with only $1212$ out of the $254254$ identified potential faults worth repairing. Following manual inspection, we infer that this happens because: (1) graph transformations across Source and Target may result in syntactic difference but be semantically equivalent. For instance, a BatchNorm node is transformed in Mul and Add nodes. Although this is localized as a potential fault by Fix-Con  the repair employed by replacing Mul and Add nodes with BatchNorm will not reduce the output label discrepancy and is therefore rejected; (2) the LH difference detected is already mitigated by the converter by introducing additional nodes in the TargetḞor example, a pads (padding) hyperparameter in a Conv node on Source is replaced by a dominating Pad to the correspondent Conv node in target, leading to the same effect.
VI-D Repair Validity
In order to verify our results, we used the ImageNetV2 [69] dataset consisting of 10K images which was built with the purpose of providing test data for ImageNet, as an independent validation source. We checked the discrepancies of the repaired Target models against Source, and verified that (1) all cases that were fully repaired using ILSVRC2017 also presented no discrepancies with ImageNetV2, and (2) partially repaired cases #$22$ and $88$ presented 0.280.280.280.28% and 0.660.660.660.66% discrepancies, lower than those observed in ILSVRC2017. The full comparison can be found in Table III. In addition, we used ImageNetV2 to verify that repaired models were giving consistent results against the dataset ground truths. Finally, we tested all $1010$ cases shown in Figure 1 presenting no discrepancies with ImageNetV2, to verify that no additional errors were detected.
VI-E Results Generalizability
Our approach is applicable to any DL framework or conversion tool. In addition, it can be applied to any DL model architecture, including non-CNN models, only requiring the user to specify the core computational layers to be used in the comparison. We chose image recognition models in our experiments as they are more tractable and need less resources than other DL architectures like transformers, as our experiments involved $1515$ model conversion cases with numerous iterations and analysis within each. All the models selected are widely used, either as-is or as the base for extension of models deployed on more complex tasks (e.g., ResNet101 extended for object detection).
In terms of the fault types supported by Fix-Con, they are based on a survey of issues encountered in discussion forums and empirical investigations (Table I). There may be other fault types not currently considered by Fix-Con. We are aware of this and we expect that Fix-Con will keep evolving to handle emerging issues and other fault categories in model conversions between DL frameworks.
TABLE III: Experiments Overview: Discrepancies between Source and Target. before and after the repair process. for ILSVRC2017 and ImageNetV2 datasets.
ILSVRC2017 ImageNetV2
# MobileNetV2 Initial Fixed Initial Fixed
1 Keras-to-Torch 83.41% 0% 74.85% 0%
2 Keras-to-TFLite 49.34% 0.33% 31.87% 0.28%
3 Torch-to-TF 68.21% 0% 53.77% 0%
4 Torch-to-Keras 47.21% 0% 26.44% 0%
5 TF-to-TFLite 9.70% 0% 6.32% 0%
6 TF-to-Keras 48.90% 0% 31.56% 0%
ResNet101
7 Keras-to-Torch 36% 0% 89% 0%
8 Keras-to-TFLite 3.12% 1.50% 1.58% 0.66%
9 Torch-to-TFLite 72.30% 0% 53.41% 0%
10 Torch-to-TF 73.90% 0% 56.15% 0%
11 TFLite-to-Keras 100.00% 0% 100.00% 0%
12 TF-to-TFLite 2.70% 2.70% 1.32% 1.32%
13 TF-to-Keras 34.85% 0% 17.59% 0%
InceptionV3
14 Torch-to-Keras 32.38% 0% 15.89% 0%
15 TF-to-TFLite 4.36% 0% 1.53% 0%
VII Conclusion
We presented Fix-Con  an automated approach for fault localization and repair during model conversion between deep learning frameworks. Fix-Con is capable of detecting and fixing six commonly encountered fault types that are input and layer-based related to pre-processing, input dimensions, layer weights, biases, hyperparameters, and computation graph. We evaluated the effectiveness of Fix-Con in fixing model conversion bugs reported by Louloudakis et al. [63] for three widely used image recognition models converted across four different deep learning frameworks. We found that Fix-Con was highly effective in repairing the erroneous model conversions reported by completely fixing or significantly improving $1414$ out of the $1515$ cases under test, repairing a total of $462462$ bugs.

### v4

@@
-Fix-Con: Automatic Fault Localization and Repair of Deep Learning Model Conversions between Frameworks
-Abstract
+FetaFix: Automatic Fault Localization and Repair of Deep Learning Model Conversions
+Abstract.
 Converting deep learning models between frameworks is a common step to maximize model compatibility across devices and leverage optimization features that may be exclusively provided in one deep learning framework.
 However, this conversion process may be riddled with bugs, making the converted models either undeployable or problematic, considerably degrading their prediction correctness.
-In this paper we propose an automated approach for fault localization and repair, Fix-Con, during model conversion between deep learning frameworks.
-Fix-Con is capable of detecting and fixing faults introduced in model input, parameters, hyperparameters, and the model graph during conversion.
-Fix-Con uses a set of fault types (mined from surveying conversion issues reported in code repositories and forums) to localize potential conversion faults in the converted target model and then repair them appropriately, e.g., replacing the parameters of the target model with those from the source model.
+In this paper, we propose an automated approach for fault localization and repair, FetaFix, during model conversion between deep learning frameworks.
+FetaFix is capable of detecting and fixing faults introduced in model input, parameters, hyperparameters, and the model graph during conversion.
+FetaFix uses a set of fault types (mined from surveying common conversion issues reported in code repositories and forums) to localize potential conversion faults in the converted target model and then repair them appropriately, e.g., replacing the parameters of the target model with those from the source model.
 This is done iteratively for every image in the dataset, comparing output label differences between the source model and the converted target model until all differences are resolved.
-We evaluate the effectiveness of Fix-Con in fixing model conversion bugs of three widely used image recognition models converted across four different deep learning frameworks.
-Overall, Fix-Con was able to fix $462462$ out of $755755$ detected conversion faults, either completely repairing or significantly improving the performance of $1414$ out of the $1515$ erroneous conversion cases.
-I Introduction
-With the widespread use of deep learning (DL) in various domains, there is an inherent need for DL models to be inter-operable across DL frameworks (such as PyTorch [67], TensorFlow (TF) [64], Keras [54]) to maximize re-usability of models across frameworks.
-Conversion of DL models between frameworks is facilitated by a plethora of automated conversion tools such as tf2onnx [44], onnx2keras [45], onnx2torch [46], MMdnn [61], among others.
-However, the conversion process may be riddled with bugs [63, 62, 43], making the converted models undeployable, perform poorly in terms of output label correctness when changing from one framework to another, run slowly, or face challenges in robust deployment [66, 52, 53].
-Louloudakis et al.
-[63] provided empirical evidence of model conversion errors and resulting output label errors in four DL frameworks.
-In this paper, we target the conversion errors they observed for fault localization and repair.
-They considered three image recognition models pre-trained on ImageNet [55] that were treated as Source models.
-They then converted each Source model to use a different DL framework, referred to as Target.
-Figure 1 from [63] shows the proportion of output label dissimilarities between (Source, Target) pairs across all images in the test dataset.
+We evaluate the effectiveness of FetaFix in fixing model conversion bugs of three widely used image recognition models converted across four different deep learning frameworks.
+Overall, FetaFix was able to fix $462462$ out of $755755$ detected conversion faults, either completely repairing or significantly improving the performance of $1414$ out of the $1515$ erroneous conversion cases.
+conference: The 29th International Conference on Evaluation and Assessment in Software Engineering; 17–20 June, 2025; Istanbul, Türkiyepublicationmonth: 06journalyear: 2025
+1. Introduction
+With the widespread use of deep learning (DL) in various domains, there is an inherent need for DL models to be inter-operable across DL frameworks (such as PyTorch (Paszke et al., 2019), TensorFlow (TF) (Martín Abadi et al., 2015), Keras (Chollet, 2015)) to maximize re-usability of models across frameworks.
+Conversion of DL models between frameworks is facilitated by a plethora of automated conversion tools such as tf2onnx (tf2, 2022), onnx2keras (onn, 2023a), onnx2torch (onn, 2023b), MMdnn (Liu et al., 2020), among others.
+However, the conversion process may be riddled with bugs (Louloudakis et al., 2023a, 2022; TF-, 2021), making the converted models undeployable, perform poorly in terms of output label correctness when changing from one framework to another, run slowly, or face challenges in robust deployment (Openja et al., 2022b; Chen et al., 2020, 2021).
+In a previous study of ours (Louloudakis et al., 2023a), we showcased empirical evidence of problems in model conversion across four well-known DL frameworks.
+In particular, we considered three image recognition models, pre-trained on ImageNet (Deng et al., 2009), that were treated as Source models.
+We then converted each Source model to use a different DL framework, referred to as Target.
+Figure 1 from (Louloudakis et al., 2023a) shows the proportion of output label dissimilarities between (Source, Target) pairs across all images in the test dataset.
 The conversion process failed to execute in $1111$ out of the $3636$ conversions, seen as empty gray cells in Figure 1.
@@
 This study underscores the error-prone nature of the conversion process, highlighting the necessity for a technique to localize and repair faults introduced by DL framework converters.
-Figure 1: Pairwise comparison of output labels between Source and converted Target models (from [63]).
-To address this problem, in this paper we propose Fix-Con, an automated approach for fault localization and repair of erroneous model conversions between DL frameworks.
-We focus on the error cases in Figure 1 with non-zero output label differences.
-The gray boxes in Figure 1, where the conversion tools crashed, have been explored by existing tools like MMdnn [61] and complements the capability we provide in diagnosing and repairing conversions with output label differences.
-Fix-Con is capable of detecting and fixing faults introduced in model layer weights, biases, hyperparameters, and the model computation graph during conversion.
+Figure 1. Pairwise comparison of output labels between Source and converted Target models (from (Louloudakis et al., 2023a)).
+To address this problem, in this paper we propose FetaFix [^1] The source code (accompanied with detailed instructions), as well a small dataset for evaluation are available at https://github.com/luludak/FetaFix, an automated approach for fault localization and repair of erroneous model conversions between DL frameworks.
+We focus on the cases in Figure 1 that did not crash, but resulted in non-zero output label differences across Source and Target.
+The gray boxes in Figure 1, where the conversion tools crashed, have been explored by existing tools like MMdnn (Liu et al., 2020) and complements the capability we provide in diagnosing and repairing conversions with output label differences.
+FetaFix is capable of detecting and fixing faults introduced in model layer weights, biases, hyperparameters, and the model computation graph during conversion.
 It compares the Source model (in one DL framework) against the converted Target model (in a different framework) using an input dataset.
-For images with label discrepancies between Source and Target, Fix-Con assesses the difference between the models with respect to their parameters.
+For images with label discrepancies between Source and Target, FetaFix assesses the difference between the models with respect to their parameters.
 Then it uses a set of strategies to mitigate conversion errors such as the replacement of Target model parameters with those from Source.
-Our approach for Fix-Con focuses on the Target model rather than the converter tool itself for a variety of reasons.
+Our approach for FetaFix focuses on the Target model rather than the converter tool itself for a variety of reasons.
 First, errors observed in conversions may be related to problematic configurations (e.g., input preprocessing) that are not related to the conversion tool.
 Second, the source code of the conversion tool may be unavailable or proprietary, and therefore not accessible for direct repair.
-Third, the conversion process might involve manual intervention or conversion (e.g., [25, 6]), which in turn can introduce errors.
-Fourth, our methodology is not tied to any specific conversion tool/version, and can be applied to any model conversion between source and target settings, including intermediate representations like ONNX [47].
+Third, the conversion process might involve manual intervention or conversion (e.g., (man, 2024b, a)), which in turn can introduce errors.
+Fourth, our methodology is not tied to any specific conversion tool/version, and can be applied to any model conversion between source and target settings, including intermediate representations like ONNX (onn, 2023c).
 It is worth noting that conversion tools are constantly evolving, thus having a localize/repair approach independent of the tool source code accessibility and evolution would prove helpful and more dependable for AI model users.
 As an aside, the fault localization and repair techniques we propose can also be adopted by conversion tool developers to address faults associated with the tools themselves.
-We evaluate the effectiveness of Fix-Con in fixing the model conversion errors reported by Louloudakis et al [63].
-Fix-Con was able to effectively localize and repair $1414$ of the $1515$ erroneous model conversions that had non-zero output label dissimilarities.
+Focusing on the non-crashing conversions from our prior work (Louloudakis et al., 2023a), we evaluated the effectiveness of FetaFix in fixing model conversion errors.
+FetaFix was able to effectively localize and repair $1414$ of the $1515$ erroneous model conversions that had non-zero output label dissimilarities.
 In summary, we make the following contributions:
-  1. 1.
+  1. (1)
     A novel method for localizing faults in DL model conversion between a range of DL frameworks.
-  2. 2.
-    An automated fault repair framework, Fix-Con, that can fix converted Target models to match output labels with that of the Source model.
-  3. 3.
-    A comprehensive empirical evaluation of the effectiveness of Fix-Con in fault localization and repair.
-II Background
-We provide a brief background on deep learning frameworks, conversions between them, the intermediate ONNX format and the compiler framework Apache TVM that we use for implementation and analysis with Fix-Con.
-II-A Deep Learning Frameworks
+  2. (2)
+    An automated fault repair framework, FetaFix, that can fix converted Target models to match output labels with that of the Source model.
+  3. (3)
+    A comprehensive empirical evaluation of the effectiveness of FetaFix in fault localization and repair.
+2. Background
+We provide a brief background on deep learning frameworks, conversions between them, the intermediate ONNX format and the compiler framework Apache TVM that we use for implementation and analysis with FetaFix.
+2.1. Deep Learning Frameworks
 Deep Learning Frameworks provide utilities to machine learning engineers and developers in order to build, train, deploy and optimize Deep Neural Networks (DNNs).
-We focus on four actively maintained and widely-used DL frameworks in our experiments: Keras [54], PyTorch [67], TensorFlow (TF) [64], and TensorFlow Lite (TFLite - a framework focused on edge devices) [64].
-II-B Framework Conversions
-DNN models are converted from one DL framework to another, for two primary reasons: (1) portability - so that a model can be deployed on devices of different computational capabilities; and (2) supported features and/or operations, such as extended optimization support or better profiling options.
+We focus on four actively maintained and widely-used DL frameworks in our experiments: Keras (Chollet, 2015), PyTorch (Paszke et al., 2019), TensorFlow (TF) (Martín Abadi et al., 2015), and TensorFlow Lite (TFLite - a framework focused on edge devices) (Martín Abadi et al., 2015).
+2.2. Framework Conversions
+DNN models are converted from one DL framework to another, for two primary reasons: (1) portability - so that a model can be deployed on devices of different computational capabilities; and (2) supported features and/or operations across the Deep Learning Acceleration Stack (Gibson et al., 2025), such as extended optimization support or better profiling options.
 The conversion process is conducted by providing a pre-trained model built on one DL framework (e.g., PyTorch).
@@
 We call this model Target.
-There is a wide variety of open-source tools for this purpose [44, 45, 46, 61], and the conversion procedure is performed automatically.
-However, the conversion process might involve multiple converters and might also require the model to be built using an intermediate representation (e.g., ONNX [47]).
-For instance, the conversion of a model from PyTorch to Keras might require a two-step conversion: (1) PyTorch to ONNX, using the native converter of PyTorch to generate ONNX; and (2), the onnx2keras [45] converter, in order to convert the ONNX intermediate representation to Keras.
-II-C Open Neural Network Exchange (ONNX)
-ONNX [47] is an open standard for machine learning interoperability that allows the representation of DNNs using a common set of operators and file formats to enable interoperability with a variety of tools, and systems.
+There is a wide variety of open-source tools for this purpose (tf2, 2022; onn, 2023a, b; Liu et al., 2020), and the conversion procedure is performed automatically.
+However, the conversion process might involve multiple converters and might also require the model to be built using an intermediate representation (e.g., ONNX (onn, 2023c)).
+For instance, the conversion of a model from PyTorch to Keras might require a two-step conversion: (1) PyTorch to ONNX, using the native converter of PyTorch to generate ONNX; and (2), the onnx2keras (onn, 2023a) converter, in order to convert the ONNX intermediate representation to Keras.
+2.3. Open Neural Network Exchange (ONNX)
+ONNX (onn, 2023c) is an open standard for machine learning interoperability that allows the representation of DNNs using a common set of operators and file formats to enable interoperability with a variety of tools, and systems.
 ONNX provides a powerful API in order to analyze model graph, hyperparameters, and parameters.
@@
 For that purpose, when we refer to Source and Target  we implicitly refer to the ONNX variants of the models under test.
-II-D Apache TVM
-Apache TVM [51] is an end-to-end machine learning compiler framework for CPUs, GPUs, and specialized accelerators.
+2.4. Apache TVM
+Apache TVM (Chen et al., 2018) is an end-to-end machine learning compiler framework for CPUs, GPUs, and specialized accelerators.
 It is actively used and supported by a wide community of developers and researchers worldwide.
 TVM allows compilation, optimization and deployment of DNNs in a variety of hardware acceleration devices.
-It also features a very powerful debugger, which allows to export metadata regarding the model inference, such as layer activations, execution times per-model layer, as well as metadata about model structure and parameters, each exported in separate metadata files, while also providing a high-level API for their analysis.
-We utilize the TVM Debugger in order to perform Layer Activation Analysis, as described in Section IV-A4.
-III Related Work
-In terms of model conversions, limited research is conducted in the literature.
+It also features a very powerful debugger, which allows exporting metadata regarding the model inference, such as layer activations, execution times per-model layer, as well as metadata about model structure and parameters, each exported in separate metadata files, while also providing a high-level API for their analysis.
+We utilize the TVM Debugger in order to perform Layer Activation Analysis, as described in Section 4.1.4.
+3. Related Work
+To achieve model conversions between DL frameworks, a variety of tools exist, such as tf2onnx (tf2, 2022), tflite2onnx (tfl, 2020), onnx2keras (onn, 2023a), onnx2torch (onn, 2023b), and native DL framework converters such as ONNX converter from PyTorch (Paszke et al., 2019) and TFLiteConverter from TFLite (Martín Abadi et al., 2015).
+However, these tools come with no guarantees, and can potentially introduce faults in the converted models.
+To this direction however, limited research is conducted in the literature.
 Openja et al.
-[65] published a study that emphasizes the challenges of the conversion process.
-Louloudakis et al.
-[63, 62] explored the effects of the computational environment parameters on image recognition models, including deep learning (DL) framework conversions.
+(Openja et al., 2022a) published a study that emphasizes the challenges of the conversion process.
+Jajal et al.
+(Jajal et al., 2024) conducted a failure analysis on ONNX model converters, concluding that converters are responsible for about 75% of the conversion defects, and that 33% of reported failures are related to semantically incorrect models.
 Liu et al.
-[61] propose MMdnn, a framework to automate the framework conversion process and mitigate a number of conversion challenges that result in conversion failures, such as unavailable operators, inconsistent tensor layouts, unsupported padding and incompatible argument types.
+(Liu et al., 2020) propose MMdnn, a framework to automate the framework conversion process and mitigate a number of conversion challenges that result in conversion failures, such as unavailable operators, inconsistent tensor layouts, unsupported padding and incompatible argument types.
 MMdnn, however, does not consider faults related to configuration, such as input preprocessing, or faults related to additional important aspects such as layer parameters and computation graph that can also result in output incompatibilities during conversion.
-This is addressed by Fix-Con.
+This is addressed by FetaFix.
 It is also worth noting that MMdnn is currently in a deprecated state, with its last update on August 2020.
 In addition, our efforts with testing MMdnn was unsuccessful as it was built with older, deprecated versions of DL frameworks (e.g., PyTorch v0.4.0, while the current version is v2.2.0).
-In an attempt to compare MMdnn with our work, we identified a potential overlap of only $1515$ potential faults out of the $462462$ we repaired with Fix-Con, based on the results reported in [61].
+We rigorously (Louloudakis et al., 2023b, 2022, a) explored the effects of the computational environment parameters on image recognition models, including deep learning (DL) framework conversions, and implemented a proof-of-concept (Louloudakis et al., 2024) concerning conversion faults related to model weights and biases, which is the predecessor of FetaFix.
+In an attempt to compare MMdnn with our work, we identified a potential overlap of only $1515$ potential faults out of the $462462$ we repaired with FetaFix, based on the results reported in (Liu et al., 2020).
 Finally, MMdnn supports conversions over a restricted set of DL frameworks (primarily CNTK, CoreML, Keras, MXNet, PyTorch, TensorFlow, Caffe, and DarkNet) that is not currently maintained or updated while our approach is generalizable to model conversions between any pair of DL frameworks.
-Automatic fault localization of DNNs is also explored in the literature DeepFault [56] focuses on a suspiciousness-oriented spectrum analysis algorithm to identify problematic parts of the model that can be responsible for faults, and also proposes a method for adversarial input generation.
-DeepLocalize [74] converts the model into an imperative representation and then performs dynamic analysis on its execution traces.
-In the context of fault localization of DL frameworks, CRADLE [68] performs an execution analysis on the model graph in order to detect faults, while LEMON [73] utilizes the metrics used by CRADLE and applies mutation testing in order to identify issues in the model.
+Automatic fault localization of DNNs is also explored in the literature.
+DeepFault (Eniser et al., 2019) focuses on a suspiciousness-oriented spectrum analysis algorithm to identify problematic model parts, and proposes a method for adversarial input generation.
+DeepLocalize (Wardat et al., 2021) converts the model into an imperative representation and then performs dynamic analysis on its execution traces.
+In the context of fault localization of DL frameworks, CRADLE (Pham et al., 2019) performs an execution analysis on the model graph in order to detect faults, while LEMON (Wang et al., 2020) utilizes the metrics used by CRADLE and applies mutation testing to identify issues in the model.
 However, none of the aforementioned approaches focus on faults encountered during model conversions.
-To achieve model conversions between DL frameworks, a variety of popular tools exist, such as tf2onnx [44], onnx2keras [45], onnx2torch [46], tflite2onnx [42], and native DL framework converters such as TFLiteConverter from TFLite [64], and ONNX converter from PyTorch [67].
-However, these tools come with no guarantees, and can potentially introduce faults in the converted models.
-IV Methodology
-Figure 2: Fault localization and repair pipeline of Fix-Con handling 6 fault types, three within Input-based category and three within Layer-based category.
-Fix-Con first takes as input the dataset, Source and Target models.
+4. Methodology
+Figure 2. Fault localization and repair pipeline of FetaFix handling 6 fault types, three within Input-based category and three within Layer-based category.
+FetaFix first takes as input the dataset, Source and Target models.
+It supports any image test dataset, requiring only the definition of a dataset path in its configuration.
 It then localizes and repairs conversion errors that resulted in output label discrepancies between Source and Target.
@@
 We then import the Source-ONNX and Target-ONNX models to the TVM compiler framework to facilitate comparison.
-Fix-Con comprises two main components: 1) A Fault Localization component that identifies locations, parameters and properties in the Target model that may be potentially faulty; and 2) a Fault Repair component with a defined set of strategies to fix the faulty locations and aspects identified through fault localization.
-The full fault localization and repair pipeline of Fix-Con is presented in Figure 2.
-IV-A Fault Localization
+FetaFix comprises two main components: 1) A Fault Localization component that identifies locations, parameters and properties in the Target model that may be potentially faulty; and 2) a Fault Repair component with a defined set of strategies to fix the faulty locations and aspects identified through fault localization.
+The full fault localization and repair pipeline of FetaFix is presented in Figure 2.
+4.1. Fault Localization
 We surveyed $4040$ posts and issues from StackOverflow, GitHub, and forums such as TensorFlow and NVIDIA discussions, where model conversion errors were raised, to understand the possible locations of errors.
-Table I shows six different conversion fault types identified in these issues that were related to 1) Preprocessing (PP), (2) Input Dimensions (ID), (3) Tensor Shape and Structure (TSS), (4) Weights & Biases (WB), (5) Layer Hyperparameters (LH), and (6) Computation Graph (CG).
-We also discovered many of these fault types examining the conversion errors reported in [63].
-TABLE I: Model conversion faults from StackOverflow, GitHub issues and other forums. Faults are classified as input-based and layer-based.
+Although not exhaustive, we believe this sample to be representative of issues faced by developers, as (1) the issues we observed were among the most popular with regard to conversions and (2) the same or very similar issues appeared across different trackers and forums.
+Table 1 shows six different conversion fault types identified in these issues that were related to 1) Preprocessing (PP), (2) Input Dimensions (ID), (3) Tensor Shape and Structure (TSS), (4) Weights & Biases (WB), (5) Layer Hyperparameters (LH), and (6) Computation Graph (CG).
+We also discovered many of these fault types examining the conversion errors reported in (Louloudakis et al., 2023a).
+Table 1. Model conversion faults from StackOverflow, GitHub issues and other forums. Faults are classified as input-based and layer-based.
 # Class Fault Sources
-1 Input $P$ [3, 19, 10, 39, 7, 30]
-2 Input ID [5, 36, 37, 32, 13, 20, 23, 26, 28]
-3 Input TSS [11, 14, 21, 27, 31, 41]
-4 Layer WB [8, 9, 12, 16, 2, 29, 34, 1, 15, 35, 33, 40]
-5 Layer LH [17, 18, 50]
-6 Layer CG [4, 38, 22, 24]
+1 Input $P$ (sta, 2023b, f, d, n; pyt, 2023; ten, 2024)
+2 Input ID (sta, 2023c; ten, 2023c; sta, 2023m, k, e; git, 2023j; sta, 2023g; git, 2023m; sta, 2023i)
+3 Input TSS (git, 2023d, f, k, n, o, t)
+4 Layer WB (git, 2023b, c, e, g; ten, 2023a; sta, 2023j, l, a; ten, 2023b; git, 2023q, p, s)
+5 Layer LH (git, 2023h, i; ACervantes, 2021)
+6 Layer CG (git, 2023a, r, l; sta, 2023h)
 As seen in Figure 2, our fault localization approach starts by localizing the first three fault types: PP, ID, and TSS.
@@
 We refer to these fault types as Input-based.
-After attempting to localize input-based fault types, our approach proceeds with the remaining three fault types in Table I that belong to the Layer-based category, as it stems from differences in layer hyperparameters, weights and biases, and differences in computation graph.
+After attempting to localize input-based fault types, our approach proceeds with the remaining three fault types in Table 1 that belong to the Layer-based category, as it stems from differences in layer hyperparameters, weights and biases, and differences in computation graph.
 Comparisons for Input-Based Fault Types
 We examine and contrast input-related settings between the Source and Target models in the fault types below.
-IV-A1 Preprocessing (PP)
+4.1.1. Preprocessing (PP)
 For some instances, differences in preprocessing settings between Source and Target can result in output label differences.
 We explore the effect of different preprocessing configurations for Target using the default preprocessing setting from the official repositories of the DL frameworks both Source and Target.
-For example, in the case of converting MobileNetV2 from PyTorch to Keras, Fix-Con examines the performance of the converted Target, using both official PyTorch and Keras preprocessing settings for MobileNetV2.
-IV-A2 Input Dimensions (ID)
+For example, in the case of converting MobileNetV2 from PyTorch to Keras, FetaFix examines the performance of the converted Target, using both official PyTorch and Keras preprocessing settings for MobileNetV2.
+4.1.2. Input Dimensions (ID)
 Islam et al.
-[60] identified that changes in layer input dimensions can affect the performance of a DNN model.
+(Islam et al., 2020) identified that changes in layer input dimensions can affect the performance of a DNN model.
 We also empirically identified that model conversion tools accept misconfigurations of Target model input dimensions without notifying the user of the mismatch between Source and Target input dimensions that may potentially result in output label discrepancies.
@@
 In our fault localization approach, we check if input dimensions between Source and Target match, and report any mismatches to the fault repair component.
-IV-A3 Tensor Shape and Structure (TSS)
+4.1.3. Tensor Shape and Structure (TSS)
 Another potential error identified by Islam et al.
-[60] relates to tensor shapes and changes in model structure.
+(Islam et al., 2020) relates to tensor shapes and changes in model structure.
 We consider cases where an erroneous tensor shape can affect the correctness of the model conversion where Source and Target models have differences in output labels.
@@
 Owing to the ambiguity in dimensions where both transpose operations result in a tensor with correct shape expected by the subsequent layers in the Target model, it can result in model conversion errors as one of them has an incorrect structure.
-To localize faults like this, Fix-Con checks if there is a difference in model input shape between Source and Target.
-For cases presenting differences, Fix-Con further examines if any of the input-related subsequent layers contains a shape transformation operation (like Transpose) and considers such scenarios as potential sources of faults, conservatively.
+To localize faults like this, FetaFix checks if there is a difference in model input shape between Source and Target.
+For cases presenting differences, FetaFix further examines if any of the input-related subsequent layers contains a shape transformation operation (like Transpose) and considers such scenarios as potential sources of faults, conservatively.
 Potential faults related to tensor structure may also happen when a layer tensor in the Target model is incorrectly transformed by the converter and is followed by a flattening or a reshaping operation that loses the original structure information, making the difference with Source model non-obvious.
@@
 For this category of faults, we examine and compare layer activations between Source and Target models.
-For layers where differences in activations are outside of an expected range of differences (as described in Section IV-A4), we mark those layers as suspicious and rank them so that we examine the most suspicious layer first for fault types CG, LH, and WB.
+For layers where differences in activations are outside of an expected range of differences (as described in Section 4.1.4), we mark those layers as suspicious and rank them so that we examine the most suspicious layer first for fault types CG, LH, and WB.
 We then iteratively proceed down this ranked list of suspicious layers for inspecting the three layer-based fault types, CG, LH and WB, discussed below.
 We start by describing our approach for producing a ranked list of suspicious layers in layer activation analysis followed by our fault localization approach for CG, LH, and WB with the ranked suspicious layers.
-An example instance of the differences between Source and Target in weights, biases, and hyperparameters that Fix-Con attempts to localize as potential faults is shown in Figure 3.
+An example instance of the differences between Source and Target in weights, biases, and hyperparameters that FetaFix attempts to localize as potential faults is shown in Figure 3.
 We find, for a core computational layer under inspection, that the weights present small differences between Source and Target.
 Additionally, the pads hyperparameter that is present in Source, is missing on Target in Figure 3.
-IV-A4 Layer Activation Analysis
+4.1.4. Layer Activation Analysis
 We compare the activations for each convolutional layer between Source and Target models, as a four-step process.
 (1) We extract two equally sized sets of images from the test dataset: one for similar images (images that produce the same top label prediction across Source and Target), and one for dissimilar images (images with different labels between Source and Target).
-We then extract the layer activation tensors of both sets by loading the Source and Target ONNX models to TVM and performing inference using the TVM Debugger [49].
+We then extract the layer activation tensors of both sets by loading the Source and Target ONNX models to TVM and performing inference using the TVM Debugger (tvm, 2023).
 The TVM debugger generates the complete model graph and parameters at compile time, while also gathering layer activations for each input.
 Using these activations and the Source ONNX model, we then utilize the LAYER_ANALYSIS function, shown in Algorithm 1.
-This function takes as input the original source ONNX model, as well as the activations for the sets of similar and dissimilar images (sim_acts, dis_acts).
+This function takes as input the original source ONNX model (used to iterate the matching layer activations for both Sourceand Target), as well as the activations for the sets of similar and dissimilar images for both Sourceand Target(contained in parameters sim_acts and diss_acts).
 It also utilizes the List module, responsible for list length check and sorting operations, and the Kruskal-Wallis parametric test (KWTest).
-(2) To identify when activation mismatches between Source and Target indicate a potential problem, we first extract the differences in activations between Source and Target, per element within each layer, for the set of similar images and store these differences in a set, sim_diff in line $88$ of Algorithm 1.
+(2) To identify when activation mismatches between Source and Target indicate a potential problem, we first extract the differences in activations between Source and Target, per element within each layer, for the set of similar images and store these differences in a set, sim_diff in lines $88$ and $99$ of Algorithm 1.
 We treat differences in activations between Source and Target for these similar images as acceptable differences and we use it to estimate the expected difference in distribution of activations for correct conversions.
-(3) Next, for the set of dissimilar images, we compute the set of dissimilar differences, diss_diff in line $1010$, per tensor element per layer.
+(3) Next, for the set of dissimilar images, we compute the set of dissimilar differences, diss_diff (lines $1111$ and $1212$), per tensor element per layer.
 We then check if the difference in activations per tensor element per layer belongs to the corresponding expected distribution of activation differences (from the previous step).
-To do this, we use the Kruskal-Wallis non-parametric statistical test (line $1111$) [57] [^1] The distribution of activation differences for correct conversion instances did not follow a normal distribution, so we chose the non-parametric Kruskal-Wallis test.
-to determine if sim_diff and diss_diff belong to the same distribution (at 5% significance level) (line $1515$).
+To do this, we use the Kruskal-Wallis non-parametric statistical test (line 13131313(et al., 2014) [^2] The distribution of activation differences for correct conversion instances did not follow a normal distribution, so we chose the non-parametric Kruskal-Wallis test.
+to determine if sim_diff and diss_diff belong to the same distribution (using a standard 5% significance level) (line $1717$).
 (4) From step 3, we obtain the tensor elements whose difference in activations for sim_diff and diss_diff did not belong to the same distribution, for each layer between Source and Target.
 These are considered as potentially problematic elements that may result in model conversion errors.
-We count the number of problematic tensor elements for each model layer (line $1616$) and rank the layers in descending order of number of problematic elements (line $1818$).
+We count the number of problematic tensor elements for each model layer (line $1818$) and rank the layers in descending order of number of problematic elements (line $2020$).
 This ranking of suspicious layers is the basis for fault localization and repair of fault types CG, LH, and WB.
@@
 However, this can be modified to consider other layer types as well.
-Algorithm 1 Layer Analysis in Fix-Con
+Algorithm 1 Layer Analysis in FetaFix
 1:List, KWTest, model, sim_acts, diss_acts
@@
 8:             for sim in sim_acts do
-9:                 sim_diff $\leftarrow$ sim_diff $\cup$ sim[element]
-10:             for diss in diss_acts do
-11:                 diss_diff $\leftarrow$ diss_diff $\cup$ diss[element]
-12:             kw_out $\leftarrow$ KWTest(sim_diff, diss_diff)
-13:             // If the null hypothesis of
-14:             // same distribution is rejected,
-15:             // then the element counts as problematic.
-16:             if kw_out.p_value <<< 0.05 then
-17:                 layer.prob_elems+++++ +
-18:    // Sort layers in descending order of problematic elements.
-19:    return sort(model.layers,
-              (l1, l2) $\Rightarrow$ l1.prob_elems -- l2.prob_elems)
-IV-A5 Weights and Biases (WB)
+9:                 sim_diff $\leftarrow$ sim_diff $\cup$
+10:                                (sim.tgt[element] - sim.src[element])
+11:             for diss in diss_acts do
+12:                 diss_diff $\leftarrow$ diss_diff $\cup$
+13:                                (diss.tgt[element] - diss.src[element])
+14:             kw_out $\leftarrow$ KWTest(sim_diff, diss_diff)
+15:             // If the null hypothesis of
+16:             // same distribution is rejected,
+17:             // then the element counts as problematic.
+18:             if kw_out.p_value <<< 0.05 then
+19:                 layer.prob_elems+++++ +
+20:    // Sort layers in descending order of problematic elements.
+21:    return sort(model.layers,
+22:                    (l1, l2) $\Rightarrow$ l1.prob_elems -- l2.prob_elems)
+4.1.5. Weights and Biases (WB)
 We extract the weight and bias tensors for each of the convolutional layers.
@@
 We output all the layers that have differences identified in weights and bias tensors between Source and Target  and pass it to the repair component.
-IV-A6 Layer Hyperparameters (LH)
+4.1.6. Layer Hyperparameters (LH)
 Incorrectly converted hyperparameters are another potential source of error.
@@
 If we find any differences, we mark them as potential faults.
-In particular, we observe $33$ types of differences: (1) a hyperparamater is present in Source but not in Target; (2) a hyperparameter is present in Target but not in Source (e.g., added by the converter tool); and (3) a hyperparameter is present both in Source and Targetbut their values are different.
-IV-A7 Computation Graph (CG)
+In particular, we observe $33$ types of differences: (1) a hyperparameter is present in Source but not in Target; (2) a hyperparameter is present in Target but not in Source (e.g., added by the converter tool); and (3) a hyperparameter is present both in Source and Targetbut their values are different.
+4.1.7. Computation Graph (CG)
 In our experiments, we found that output label discrepancies can sometimes arise from differences in model graph between Source and Target models.
@@
 We repeat this for any other suspicious layers in the ranked list.
-Figure 3: Indicative example of differences in layer weights and hyperparameters, introduced in the model conversion process.
-IV-B Fault Repair
+Figure 3. Indicative example of differences in layer weights and hyperparameters, introduced in the model conversion process.
+4.2. Fault Repair
 We implemented a number of strategies to repair the potential faults identified by the fault localization component.
@@
 Input-based Repair Strategies
-IV-B1 Preprocessing Differences (PP) Repair
+4.2.1. Preprocessing Differences (PP) Repair
 We experiment running the Target model with both the Target and Source preprocessing settings to check if the label discrepancies are resolved.
-IV-B2 Input Dimensions (ID) Repair
-Following fault localization, if a difference is detected between Source and Target input dimensions, Fix-Con automatically fixes Target to match with the input dimensions from Source.
+4.2.2. Input Dimensions (ID) Repair
+Following fault localization, if a difference is detected between Source and Target input dimensions, FetaFix automatically fixes Target to match with the input dimensions from Source.
 We observed four erroneous model conversions in our experiment with incorrect input dimensions, with two of them requiring further actions in model nodes for tensor shape and structure (discussed below).
-IV-B3 Tensor Shape and Structure (TSS) Repair
-Following fault localization, as described in Section IV-A3, if the input is different between Target and Source in a problematic conversion, and input is succeeded by a Transpose node, then Fix-Con attempts to apply a fix by removing that node, in addition to adjusting the Target input to match Source.
-Furthermore, if Fix-Con detects any nodes that are related to input processing and normalization (e.g., InceptionV3 architecture contains Gather, Unsqueeze, Add, and Mul nodes before the added Transpose), then Fix-Con adjusts the attributes of these nodes in order to be compatible to the Source input dimension.
-In cases where Fix-Con detects the presence of Flatten or Reshape nodes that alter the shape of tensors, then it attempts to repair this by inserting a Transpose node right before Flatten or Reshape to match tensor shape with Source.
+4.2.3. Tensor Shape and Structure (TSS) Repair
+Following fault localization, as described in Section 4.1.3, if the input is different between Target and Source in a problematic conversion, and input is succeeded by a Transpose node, then FetaFix attempts to apply a fix by removing that node, in addition to adjusting the Target input to match Source.
+Furthermore, if FetaFix detects any nodes that are related to input processing and normalization (e.g., InceptionV3 architecture contains Gather, Unsqueeze, Add, and Mul nodes before the added Transpose), then FetaFix adjusts the attributes of these nodes in order to be compatible to the Source input dimension.
+In cases where FetaFix detects the presence of Flatten or Reshape nodes that alter the shape of tensors, then it attempts to repair this by inserting a Transpose node right before Flatten or Reshape to match tensor shape with Source.
 Layer-based Repair Strategies
-IV-B4 Weights and Biases (WB) Repair
-If a mismatch is found between any weight or bias tensor across model layers by the fault localization component, then Fix-Con replaces the values from the corresponding tensor in Source to Target.
-As mentioned earlier, Fix-Con performs the repair starting from the most suspicious layer in the ranked list.
-IV-B5 Layer Hyperparameters (LH) Repair
-Upon marking a layer hyperparamter case as potentially problematic, as described in Section IV-A6, Fix-Con attempts to fix Target by applying the corresponding setting of Source.
-In particular, if (1) a node is present in Source but is missing from the computation graph of Target, Fix-Con will update the corresponding node including the Source hyperparameter to Target.
-If (2) a hyperparameter is added in Target but is missing from Source, Fix-Con will remove it from the target node.
-And finally, if (3) there is a value mismatch between Source and Target for the hyperparameter, Fix-Con will update Target in accordance to the source value.
-IV-B6 Computation Graph (CG) Repair
-In accordance to Section IV-A7, if a difference in subgraphs is detected between Source and Target, Fix-Con attempts to repair it by replacing the Target subgraph nodes with Source subgraph nodes, preserving all the node properties and ensuring all weights, biases and hyperparameters are correctly updated in the Target graph.
-Regarding the order in which our repair strategies are applied, we first applied the ones that were input-based, namely PP, ID, and TSS.
+4.2.4. Weights and Biases (WB) Repair
+If a mismatch is found between any weight or bias tensor across model layers by the fault localization component, then FetaFix replaces the values from the corresponding tensor in Source to Target.
+As mentioned earlier, FetaFix performs the repair starting from the most suspicious layer in the ranked list.
+4.2.5. Layer Hyperparameters (LH) Repair
+Upon marking a layer hyperparameter case as potentially problematic, as described in Section 4.1.6, FetaFix attempts to fix Target by applying the corresponding setting of Source.
+In particular, if (1) a node is present in Source but is missing from the computation graph of Target, FetaFix will update the corresponding node including the Source hyperparameter to Target.
+If (2) a hyperparameter is added in Target but is missing from Source, FetaFix will remove it from the target node.
+And finally, if (3) there is a value mismatch between Source and Target for the hyperparameter, FetaFix will update Target in accordance to the source value.
+4.2.6. Computation Graph (CG) Repair
+In accordance to Section 4.1.7, if a difference in subgraphs is detected between Source and Target, FetaFix attempts to repair it by replacing the Target subgraph nodes with Source subgraph nodes, preserving all the node properties and ensuring all weights, biases and hyperparameters are correctly updated in the Target graph.
+Regarding the order in which our repair strategies are applied, we first applied the ones that were input-based, namely PP, ID, and TSS, considering that issues introduced at this early processing stage can lead to problematic behavior accumulating and spreading throughout the model.
 We do PP first as changes to preprocessing may fix errors observed downstream.
@@
 Note that some graph transformations may also be related to hyperparameter changes, e.g., removing a Target Pad node to adapt to Source structure requires the addition of a padding hyperparameter to the next convolutional layer.
-We implemented Fix-Con in Python.
+We implemented FetaFix in Python.
 We used Apache TVM Debugger for layer activations analysis.
@@
 In order to implement the repair strategies, we used the ONNX API to modify Target.
-V Experiments
-We consider three widely used image recognition models of various sizes: MobileNetV2 [71], ResNet101V2 [58, 59], and InceptionV3 [72].
-We use models pre-trained on ImageNet [55] using native model definitions and pre-trained parameters/weights sourced from four different DL frameworks’ repositories: Keras [54], PyTorch [67], TensorFlow(TF) [64], and TFLite [64].
-For our experiments, we used two datasets: the ILSVRC2017 [70] object detection test dataset consisting of $55005500$ images, and the ImageNetV2 dataset [69] consisting of 10,0001000010,00010 , 000 images.
+5. Experiments
+We consider three widely used image recognition models of various sizes: MobileNetV2 (Sandler et al., 2018), ResNet101V2 (He et al., 2015, 2016), and InceptionV3 (Szegedy et al., 2015).
+We use models pre-trained on ImageNet (Deng et al., 2009) using native model definitions and pre-trained parameters/weights sourced from four different DL frameworks’ repositories: Keras (Chollet, 2015), PyTorch (Paszke et al., 2019), TensorFlow(TF) (Martín Abadi et al., 2015), and TFLite (Martín Abadi et al., 2015).
+For our experiments, we used two datasets: the ILSVRC2017 (Russakovsky et al., 2015) object detection test dataset consisting of $55005500$ images, and the ImageNetV2 dataset (Recht et al., 2019) consisting of 10,0001000010,00010 , 000 images.
 We used the ILSVRC2017 dataset for detecting and repairing errors in model conversion.
-We then assessed the correctness of the repaired models on ImageNetV2, utilizing it as an independent dataset.
-We used the problematic conversion cases identified in [63] as our candidates for fault localization and repair.
+We then assessed the correctness of the repaired models on ImageNetV2, utilizing it as an independent dataset, to verify our results and avoid dataset overfitting.
+We used the problematic conversion cases identified in (Louloudakis et al., 2023a) as our candidates for fault localization and repair.
+Both datasets contain real, non-synthetic samples and they are widely utilized by developers and researchers, while remaining tractable in size.
 We focused on cases where the conversion process introduced label discrepancies between Source and Target, and there were $1515$ such erroneous conversions, as seen in Figure 1.
+Overall, we selected models in the domain of computer vision for the following reasons: (1) To test our repair approach against the conversion faults we detected in the models we examined in our previous work (Louloudakis et al., 2023a); and (2) since our experiments involved 15 model conversion cases with numerous iterations and analysis within each, we chose image recognition models as they are more tractable and need less resources than other DL architectures, like transformers, for our experiments.
+However, our approach can generalize easily to any DL framework or conversion tool.
+It can be applied to any DL model architecture, including non-CNN models, only requiring the user to specify the core computational layers to be used in the comparison.
 We evaluate the following research questions in our experiments:
-RQ1:
-How effective is Fix-Con at localizing faults in the $1515$ erroneous model conversions?
-RQ2:
-How effective is Fix-Con at repairing the faults in the $1515$ erroneous model conversions? We evaluate which of the repair strategies are used most frequently in model fixes.
-To check the robustness of Fix-Con with respect to devices, we ran our experiments on two devices of varying capabilities, an Intel-based server featuring a high-end Nvidia Tesla K40c (GK11BGL) GPU, and a Laptop featuring a low-end Intel(R) GEN9 HD Graphics NEO.
+RQ1::
+How effective is FetaFix at localizing faults in the $1515$ erroneous model conversions?
+RQ2::
+How effective is FetaFix at repairing the faults in the $1515$ erroneous model conversions? We evaluate which of the repair strategies are used most frequently in model fixes.
+To check the robustness of FetaFix with respect to devices, we ran our experiments on two devices of varying capabilities, an Intel-based server featuring a high-end Nvidia Tesla K40c (GK11BGL) GPU, and a Laptop featuring a low-end Intel(R) GEN9 HD Graphics NEO.
 Our experiment results remained the same on both devices.
-V-A Fault Analysis Setup
+5.1. Fault Analysis Setup
 Within layer activation analysis, discussed in fault localization for layer-based types, we determine the expected distribution of differences in activations per layer between Source and Target for images where the output label is the same using a random sample of up to $100100$ images that produced similar results.
@@
 We repeat this process $33$ times with three different random samples of similar images.
-V-B Fault Repair Setup
-Once we extracted the order of layers from the fault analysis, we utilized the fault repair component of Fix-Con, applying the strategies in a greedy approach, as described in Section IV-B.
+5.2. Fault Repair Setup
+Once we extracted the order of layers from the fault analysis, we utilized the fault repair component of FetaFix, applying the strategies in a greedy approach, as described in Section 4.2.
 We set a time limit of 2 hours for each conversion case.
@@
 These settings can be customized as needed in our tool.
-VI Results
-Overall, Fix-Con was able to localize $755755$ potential faults in the structure and properties between Source and Target variants across all $1515$ erroneous conversion cases.
-An overview of the potential faults distribution can be found in Table II.
-From these $755755$ cases Fix-Con identified as bugs, it repaired $462462$ of them — completely repairing $1212$ out of $1515$ problematic model conversions, while significantly improving the performance of $22$ more, with the most distinctive case being improved from 48.948.948.948.9% output label discrepancy to 0.330.330.330.33%.
-Our system was unable to reduce the percentage further, as none of our strategies affected the model beyond this point.
-VI-A Fault Localization
+6. Results
+[…]
+6.1. Fault Localization
 The faults localized were distributed across all six types within input-based and layer-based fault categories.
 Of the $755755$ faults localized, $462462$ of them (61.161.161.161.1% of total cases) resulted in reducing output label discrepancy between Source and Target when repaired.
-The remaining localized faults did not result in any improvement after repair and are treated as false positives detected by the fault localization component.
-We demonstrate the distribution of faults detected and fixed across the different types in Table II and discuss each type in the context of fault localization next.
-TABLE II: Number of faults localized versus those repaired (#Localisations/#Repairs) per model conversion case.
+For the remaining localized faults, FetaFix attempted to repair them by applying the respective repair strategy, but did not result in any improvement.
+Furthermore, they are considered as false positives detected by the fault localization component and are automatically discarded from the repair process.
+We demonstrate the distribution of faults detected and fixed across the different types in Table 2 and discuss each type in the context of fault localization next.
+Overall, the faults we detected are new, as we detected no reported issues with relation to them in the respective converter repositories.
+We aim reporting the issues to the respective developers of the conversion tools.
+Table 2. Number of faults localized versus repaired (#Localisations/#Repairs) per model conversion case.
 #Localisations/#Repairs
@@
 All models presented severe performance degradation in output correctness if no preprocessing option was selected.
-Input Dimensions (ID): Fix-Con identified four model cases where the input dimensions were different between Source and Target.
+Input Dimensions (ID): FetaFix identified four model cases where the input dimensions were different between Source and Target.
 We identified that input dimension errors had two primary causes: (1) erroneous configuration from the user, which neither the model converters nor the model itself raised as warnings or errors; (2) implicit changes in model input due to the converter implementation.
-We observed two cases, $99$ and $1313$ in Table II, where ID fault is a result of the first cause.
+We observed two cases, $99$ and $1313$ in Table 2, where ID fault is a result of the first cause.
 Regarding the second cause, we detected two cases where the converter changed the model input.
 In addition to ID faults, we also detected differences in tensor shape and structure that we discuss as part of the next category.
-Tensor Shape and Structure (TSS): Fix-Con detected three cases with potential faults in the TSS category.
-The model converter introduced a Transpose node in these two cases to address the difference in ID so that subsequent nodes use inputs with the correct dimension.
-However, in these two cases getting to the correct input dimension with the Transpose node was not adequate.
-This occurred due to similar sized dimensions within tensors, like the third and fourth dimension within an example tensor with shape (1, 3, 224, 224), where it is possible to incorrectly transpose dimensions of the same size that result in a different structure from what is intended.
+Tensor Shape and Structure (TSS): FetaFix detected three cases with potential faults in the TSS category.
+The model converter introduced a Transpose node in these two cases to address the difference in ID, so that subsequent nodes use inputs with the correct dimension.
+However, in these two cases achieving the correct input dimension by using the Transpose node was not adequate.
+This occurred due to similarly sized dimensions within tensors, like the third and fourth dimension within an example tensor with shape (1, 3, 224, 224), where it is possible to incorrectly transpose dimensions of the same size that result in a different structure from what is intended.
 In addition, in some model architectures (e.g., InceptionV3) the Transpose node is placed after a series of nodes that normalize data based on a specific input dimension.
 For instance, in InceptionV3, the Transpose node is placed after a series of Gather, Unsqueeze, Mul, and Add nodes.
-Fix-Con detects these differences across model graph and reports them during fault localization.
-Weights and Biases (WB): We observed $438438$ cases of differences in tensor values between weights and biases across Source and Target, with ResNet101 converted from Keras to TFLite containing $177177$ of them.
-Layer Hyperparameters (LH): Fix-Con detected $254254$ potential faults across $1212$ model conversion cases, with most of them concerning padding hyperparameters.
-Fix-Con also identified some cases where a potential hyperparameter fault was associated with a potential computation graph (CG) fault - for instance, specific convolutional layers presented potential faults related with padding hyperparameters and Pad nodes that directly preceded these convolutional layers.
-Computation Graph (CG): Our system identified $4747$ cases across $44$ conversion cases in which parts of the model computation graph presented discrepancies across Source and Target variants.
+FetaFix detects these differences across model graph and reports them during fault localization.
+Weights and Biases (WB): We observed $438438$ cases of tensor value differences in weights and biases between Source and Target, with ResNet101 converted from Keras to TFLite containing $177177$ of them.
+Layer Hyperparameters (LH): FetaFix detected $254254$ potential faults across $1212$ model conversion cases, with most of them concerning padding hyperparameters.
+FetaFix also identified some cases where a potential hyperparameter fault was associated with a potential computation graph (CG) fault - for instance, specific convolutional layers presented potential faults related with padding hyperparameters and Pad nodes that directly preceded these convolutional layers.
+Computation Graph (CG): FetaFix identified $4747$ cases across $44$ conversion cases in which parts of the model computation graph presented discrepancies across Source and Target variants.
 All the cases observed were associated primarily with three operations: (1) setting a padding around a tensor, by utilizing Pad nodes; (2) performing batch normalization to a tensor by applying an element-by-element constant addition, and scaling using a BatchNorm node; and (3) redundant output nodes erroneously placed by the converter.
-From those $4747$ cases, Fix-Con successfully identified $44$ of those cases as actual bugs and repaired them, playing a crucial role for the repair of cases #$1111$ and $1313$.
-VI-B Fault Repair
+From those $4747$ cases, FetaFix successfully identified $44$ of them as actual bugs and repaired them, playing a crucial role for the repair of cases #$1111$ and $1313$.
+6.2. Fault Repair
 […]
-PP: As shown in Table III, Fix-Con was able to repair cases #$11$, $33$, $66$, $77$, $99$ and $1010$, by determining and setting the correct preprocessing configuration.
+PP: As shown in Table 3, FetaFix was able to repair cases #$11$, $33$, $66$, $77$, $99$ and $1010$, by determining and setting the correct preprocessing configuration.
 Given that the preprocessing is not a standard procedure across models, we identified that for these cases using the Source preprocessing instead of the one officially provided for the Target library fixed the result completely.
@@
 The additional steps are explained in more detail in the categories below.
-ID: Among the $44$ faults localized as ID, Fix-Con repaired one case completely by solely adjusting the input dimensions of Target to match Source.
+ID: Among the $44$ faults localized as ID, FetaFix repaired one case completely by solely adjusting the input dimensions of Target to match Source.
 For the remaining three cases, additional actions had to be performed, involving repair tensor shape and structure, as well as computation graph repair, as described in the next categories.
-TSS: Fix-Con was able to repair three cases related to problematic tensor shape and structure.
+TSS: FetaFix was able to repair three cases related to problematic tensor shape and structure.
 In particular, cases $44$, $1313$, and $1414$ were related to problematic layers right after input.
@@
 However, the transformation was done improperly, leading to severe performance degradation of up to 68.2168.2168.2168.21%.
-We identified all cases to be related to the onnxmltools [48] converter (part of the native ONNX codebase), when converting Keras to ONNX.
-Fix-Con automatically identified the presence of a redundant Transpose node following the input node, and repaired it by (1) adjusting the Target model input dimensions so that they correspond to those of Source, (2) altering its tensor transposition order (permutation) property to align with the new input dimensions, and (3) updating the properties of nodes connected to Transpose to ensure proper tensor handling, reflecting the change in Transpose.
-This modification in particular was vital to case #$1414$ where Fix-Con detected additional nodes (Gather, Unsqueeze, Add and Mul) preceding Transpose that apply input normalization.
-As part of the repair process, Fix-Con updated the axes properties of Gather and Unsqueeze nodes to align with the updated input dimension and the Transpose permutation in Target.
+We identified all cases to be related to the onnxmltools (onn, 2023d) converter (part of the native ONNX codebase), when converting Keras to ONNX.
+FetaFix automatically identified the presence of a redundant Transpose node following the input node, and repaired it by (1) adjusting the Target model input dimensions so that they correspond to those of Source, (2) altering its tensor transposition order (permutation) property to align with the new input dimensions, and (3) updating the properties of nodes connected to Transpose to ensure proper tensor handling, reflecting the change in Transpose.
+This modification in particular was vital to case #$1414$ where FetaFix detected additional nodes (Gather, Unsqueeze, Add and Mul) preceding Transpose that apply input normalization.
+As part of the repair process, FetaFix updated the axes properties of Gather and Unsqueeze nodes to align with the updated input dimension and the Transpose permutation in Target.
 WB: Cases #$22$, $55$, $88$, and $1515$ were related to parameter precision discrepancies between Source and Target.
 Although the discrepancies were negligible, the model behavior deviated across the cases from 3.123.123.123.12-9.79.79.79.7%.
-Fix-Con performed multiple iterations and managed to completely fix cases #$55$ and $1515$ while also significantly improving cases $22$ and $88$ by replacing the weights of Target with those of Sourcewhere differences were detected.
+FetaFix performed multiple iterations and managed to completely fix cases #$55$ and $1515$ while also significantly improving cases $22$ and $88$ by replacing the weights of Target with those of Source where differences were detected.
 We identified all cases to be related to the behavior and the setup of TFLiteConverter.
 The discrepancy after each repair step is depicted in Figure 4.
-At each repair step, Fix-Con adjusts the weights & biases of a suspicious layer in Target with those from Source, in the order determined by the Layer Activation Analysis (Section IV-A4), and evaluates the effects of the change by comparing the predictions of Source and Target.
-LH: From the $254254$ potential faults Fix-Con localized, only $1212$ of them were actively affecting the model.
+At each repair step, FetaFix adjusts the weights & biases of a suspicious layer in Target with those from Source, in the order determined by the Layer Activation Analysis (Section 4.1.4), and evaluates the effects of the change by comparing the predictions of Source and Target.
+LH: From the $254254$ potential faults FetaFix localized, only $1212$ of them were actively affecting the model.
 These hyperparameters were associated with the pads hyperparameter of Conv and MaxPool nodes (cases #$1111$ & $1313$), as well as with the axes of Gather and Unsqueeze nodes responsible for input processing (case #$1414$).
 Repairing Target entailed adjusting their values with those from Source.
-CG: Fix-Con identified $4747$ cases of graph-related potential faults.
+CG: FetaFix identified $4747$ cases of graph-related potential faults.
 Of these, $44$ cases resulted in fault repairs that improved model performance (reduced output label discrepancy).
@@
 In addition, three problematic padding nodes were found to affect the behavior of the models.
-Fix-Con identified such faults and completely repaired the models by (1) removing any additional output nodes added to the model by the converter, and (2) repairing problematic padding nodes.
+FetaFix identified such faults and completely repaired the models by (1) removing any additional output nodes added to the model by the converter, and (2) repairing problematic padding nodes.
 For all the false positive cases observed in LH and CG, we performed manual inspection and observed that the converters performed semantically equivalent computation graph conversions.
@@
 Also, BatchNorm nodes of Source were replaced by the converter with Mul and Add nodes in Target  effectively applying the normalization in a two-step semantically equivalent manner.
-Fix-Con attempted to repair these scenarios by replacing them with the corresponding Source subgraph but these fixes had no effect on the number of output label discrepancies encountered.
+FetaFix attempted to repair these scenarios by replacing them with the corresponding Source subgraph but these fixes had no effect on the number of output label discrepancies encountered.
 In terms of the number of iterations needed during the repair process, input-based strategies are only applied once.
@@
 Four erroneous model conversions that required multiple repair iterations related to WB strategy repair are presented in Figure 4, ranging from $22$ to $77$ repair cycles, using the ILSVRC2017 dataset.
-Figure 4: Model conversion cases that required an iterative weights and biases repair strategy with the percentage output label dissimilarity shown after each repair cycle (following preprocessing repair).
-VI-C Fault Localization/Repair Effectiveness
+Figure 4.
+Model conversion cases that required an iterative weights and biases repair strategy with the percentage output label dissimilarity shown after each repair cycle (following preprocessing repair).
+6.3. Fault Localization/Repair Effectiveness
 We consider effectiveness as the number of localized faults that could be repaired to improve Target prediction accuracy towards matching this of Source.
-We go through each fault category to determine the effectiveness of the fault localization component, as seen in the Total row in Table II.
-We find that Fix-Con had an 100100100100% efficacy rate (all potential fault occurrences detected could be repaired) for preprocessing (PP: 9/9999/99 / 9), input dimensions (ID: 4/4444/44 / 4) and tensor structure (TSS: 2/2222/22 / 2), as well as 98.198.198.198.1% efficacy for weights and biases (WB: 430/438430438430/438430 / 438).
+We go through each fault category to determine the effectiveness of the fault localization component, as seen in the Total row in Table 2.
+We find that FetaFix had an 100100100100% efficacy rate (all potential fault occurrences detected could be repaired) for preprocessing (PP: 9/9999/99 / 9), input dimensions (ID: 4/4444/44 / 4) and tensor structure (TSS: 2/2222/22 / 2), as well as 98.198.198.198.1% efficacy for weights and biases (WB: 430/438430438430/438430 / 438).
 The weights and biases category had the most faults detected and repaired.
-Fix-Con had much less success on cases related to the computation graph (CG), with $44$ of the $4747$ potential faults detected being repaired (8.58.58.58.5%) while other cases when repaired did not result in reduction of output label discrepancy.
+FetaFix had much less success on cases related to the computation graph (CG), with $44$ of the $4747$ potential faults detected being repaired (8.58.58.58.5%) while other cases when repaired did not result in reduction of output label discrepancy.
 The CG fault repairs, although few in number, were instrumental in completely fixing conversion cases #$1111$ and $1313$.
@@
 For instance, a BatchNorm node is transformed in Mul and Add nodes.
-Although this is localized as a potential fault by Fix-Con  the repair employed by replacing Mul and Add nodes with BatchNorm will not reduce the output label discrepancy and is therefore rejected; (2) the LH difference detected is already mitigated by the converter by introducing additional nodes in the TargetḞor example, a pads (padding) hyperparameter in a Conv node on Source is replaced by a dominating Pad to the correspondent Conv node in target, leading to the same effect.
-VI-D Repair Validity
-In order to verify our results, we used the ImageNetV2 [69] dataset consisting of 10K images which was built with the purpose of providing test data for ImageNet, as an independent validation source.
+Although this is localized as a potential fault by FetaFix  the repair employed by replacing Mul and Add nodes with BatchNorm will not reduce the output label discrepancy and is therefore rejected; (2) the LH difference detected is already mitigated by the converter by introducing additional nodes in the Target.
+For example, a pads (padding) hyperparameter in a Conv node on Source is replaced by a dominating Pad to the correspondent Conv node in target, leading to the same effect.
+In terms of the fault analysis for layer-based types, we set FetaFix to perform fault localization with and without fault analysis, and we compared the outcomes.
+We concluded that FetaFix was able to repair the models in a similar manner with the results presented without the fault analysis, however enabling it applied fewer fault repairs, more effectively avoiding false positive cases where the applied repair was ineffective (up to 3.6% reduction).
+We also found the layer ranking assisted towards applying smaller repair steps, with the dissimilarity percentage decreasing more gradually with each repair cycle, making it easy to monitor and control the fault process.
+6.4. Repair Validity
+In order to verify our results, we used the ImageNetV2 (Recht et al., 2019) dataset consisting of 10K images which was built with the purpose of providing test data for ImageNet, as an independent validation source.
 We checked the discrepancies of the repaired Target models against Source, and verified that (1) all cases that were fully repaired using ILSVRC2017 also presented no discrepancies with ImageNetV2, and (2) partially repaired cases #$22$ and $88$ presented 0.280.280.280.28% and 0.660.660.660.66% discrepancies, lower than those observed in ILSVRC2017.
-The full comparison can be found in Table III.
+The full comparison can be found in Table 3.
 In addition, we used ImageNetV2 to verify that repaired models were giving consistent results against the dataset ground truths.
 Finally, we tested all $1010$ cases shown in Figure 1 presenting no discrepancies with ImageNetV2, to verify that no additional errors were detected.
-VI-E Results Generalizability
-Our approach is applicable to any DL framework or conversion tool.
-In addition, it can be applied to any DL model architecture, including non-CNN models, only requiring the user to specify the core computational layers to be used in the comparison.
-We chose image recognition models in our experiments as they are more tractable and need less resources than other DL architectures like transformers, as our experiments involved $1515$ model conversion cases with numerous iterations and analysis within each.
+In detail, we verified in our experiments that the results of each repair cycle do not misclassify previously correct results, by testing all repaired models on ImageNetV2, checking the classification in two ways - against ground truth for the dataset inputs, and by comparing the results between source and target models.
+We did not observe any misclassification cases of correct Target model predictions prior to the repair process across the entirety of our experiment set.
+6.5. Root Causes of Faults
+We consider that many of the errors detected might have common root causes as part of the conversion process.
+Upon further manual examination, we observed that (1) all faults related with preprocessing could be repaired when the proper setting was selected, (2) errors related to Weights & Biases (WB) were related to an implicit action caused by tf2onnx converter, and (3) issues related to Input Dimensions (ID), Tensor Shape & Structure (TSS), Layer Hyperparameters (LH) and Computational Graph (CG) were attributed to problematic handling of the layers by the conversion tools (e.g., both torch.onnx and onnx2keras converters mishandled model inputs).
+We aim to report such issues to the development teams of each converter tool.
+6.6. Results Generalizability
+We expect our approach to be applicable to any DL frameworks containing core computational layers and their conversion tools - including non-CNN models, requiring from the user to specify the core computational layers to be used in the comparison.
+In addition, the user must define a comparator for the evaluation of the model type under test (e.g., BLEU (Papineni et al., 2002) for text generation models).
+Providing proof of FetaFix effectiveness in this scenario, is subject to future work.
+In terms of our experiments, we chose image recognition models in our experiments as they are more tractable and need less resources than other DL architectures like transformers, as our experiments involved $1515$ model conversion cases with numerous iterations and analysis within each.
 All the models selected are widely used, either as-is or as the base for extension of models deployed on more complex tasks (e.g., ResNet101 extended for object detection).
-In terms of the fault types supported by Fix-Con, they are based on a survey of issues encountered in discussion forums and empirical investigations (Table I).
-There may be other fault types not currently considered by Fix-Con.
-We are aware of this and we expect that Fix-Con will keep evolving to handle emerging issues and other fault categories in model conversions between DL frameworks.
-TABLE III: Experiments Overview: Discrepancies between Source and Target. before and after the repair process. for ILSVRC2017 and ImageNetV2 datasets.
+Our current implementation is bassed on ONNX, as it is a widely-used format, providing a comprehensive API for model definition and modification.
+However, our methodology can be easily extended to other representations.
+In terms of the fault types supported by FetaFix, they are based on a survey of issues encountered in discussion forums and empirical investigations (Table 1).
+There may be other fault types not currently considered by FetaFix.
+We are aware of this and we expect that FetaFix will keep evolving to handle emerging issues and other fault categories in model conversions between DL frameworks.
+Table 3. Experiments Overview: Discrepancies between Source and Target. before and after the repair process. for ILSVRC2017 and ImageNetV2 datasets.
 ILSVRC2017 ImageNetV2
@@
 15 TF-to-TFLite 4.36% 0% 1.53% 0%
-VII Conclusion
-We presented Fix-Con  an automated approach for fault localization and repair during model conversion between deep learning frameworks.
-Fix-Con is capable of detecting and fixing six commonly encountered fault types that are input and layer-based related to pre-processing, input dimensions, layer weights, biases, hyperparameters, and computation graph.
-We evaluated the effectiveness of Fix-Con in fixing model conversion bugs reported by Louloudakis et al.
-[63] for three widely used image recognition models converted across four different deep learning frameworks.
-We found that Fix-Con was highly effective in repairing the erroneous model conversions reported by completely fixing or significantly improving $1414$ out of the $1515$ cases under test, repairing a total of $462462$ bugs.
+7. Challenges & Limitations
+Upon FetaFix development, we encountered a number of challenges and limitation, which we describe in this section.
+Node Matching Across Source & Target In order to perform the fault localization and repair process, FetaFix had to correctly match the nodes across Source and Target.
+However, the node matching was not one-to-one, due to (1) changes in graph structure applied by converter tools, (2) modifications in shape and internal structure of model nodes and (3) alterations in layer names between the model variants.
+To mitigate this problem, we introduced a solid set of requirements that had to be fulfilled for two nodes to be marked as the same.
+In detail, we required that two model nodes were (1) of the same type, (2) of the same shape, and (3) they shared a number of parameters on which values could only deviate by a small margin of difference (absolute value of $0.20.2$, obtained following step by step decrease to the point where the system failed to match the same nodes).
+For the moment however, we match only core computational layers (e.g., convolutions), considering that their number is guaranteed to remain unaffected by the conversion process due to model design.
+We also limit the matching algorithm to nodes containing weights and/or biases, as the presence of such properties increased significantly the matching ability of FetaFix, to the extent of more than 95% across Source and Target.
+We aim to improve our system to this direction by performing matching considering the presence of the same hyperparameters across nodes between Source and Target.
+ONNX Opset Representation Limitations: For case #$1212$, we were unable to compile both Source and Target to the same opset, despite our efforts, due to ONNX operation incompatibilities.
+For that reason, we generated Source in opset=11 and Target in opset=13.
+However, the two opsets had completely different graph representations, and FetaFix was unable to correctly match the layers across model variants and perform fault localization and repair.
+Computationally Intensive Fault Analysis: As the fault analysis we conducted performed per-element comparison for a number of inputs across each layer, it resulted in a vast amount of computations.
+This limited our tool from performing it to the full validation dataset, for reasons of tractability.
+For that matter, we selected a random sample of up to 100 images and we conducted our analysis on it, and repeated this process 3 times to solidify our results.
+We also accelerated the process by enabling parallel computation for the analysis, using Joblib python library (job, 2024).
+8. Conclusion
+We presented FetaFix, an automated approach for fault localization and repair during model conversion between deep learning frameworks.
+FetaFix is capable of detecting and fixing six commonly encountered fault types that are input and layer-based related to pre-processing, input dimensions, layer weights, biases, hyperparameters, and computation graph.
+We evaluated the effectiveness of FetaFix in fixing previously reported model conversion bugs (Louloudakis et al., 2023a) for three widely used image recognition models converted across four different deep learning frameworks.
+We found that FetaFix was highly effective in repairing the erroneous model conversions reported by completely fixing or significantly improving $1414$ out of the $1515$ cases under test, repairing a total of $462462$ bugs.
+Acknowledgements.
+The research has been funded by the Royal Society Industry Fellowship ”AutoTest: Testing Autonomous Vehicle Perception Safety on Hardware Accelerators” and Huawei Edinburgh Joint Lab Project ”RobustCheck: Testing Robustness of Compiler Optimisations and Deep Learning Frameworks”.